### v6

# Identifying Weight-Variant Latent Causal Models

Abstract
The task of causal representation learning aims to uncover latent higher-level causal representations that affect lower-level observations. Identifying true latent causal representations from observed data, while allowing instantaneous causal relations among latent variables, remains a challenge, however. To this end, we start with the analysis of three intrinsic indeterminacies in identifying latent space from observations: transitivity, permutation indeterminacy, and scaling indeterminacy. We find that transitivity acts as a key role in impeding the identifiability of latent causal representations. To address the unidentifiable issue due to transitivity, we introduce a novel identifiability condition where the underlying latent causal model satisfies a linear-Gaussian model, in which the causal coefficients and the distribution of Gaussian noise are modulated by an additional observed variable. Under some mild assumptions, we can show that the latent causal representations can be identified up to trivial permutation and scaling. Furthermore, based on this theoretical result, we propose a novel method, termed Structural caUsAl Variational autoEncoder (SuaVE), which directly learns latent causal representations and causal relationships among them, together with the mapping from the latent causal variables to the observed ones. Experimental results on synthetic and real data demonstrate the identifiability and consistency results and the efficacy of SuaVE in learning latent causal representations.
Code is available at : https://sites.google.com/view/yuhangliu/projects
Keywords: Causal Representations, Latent Causal Models, Identifiability
1 Introduction
While there is no universal formal definition, one widely accepted feature of disentangled representations (Bengio et al., 2013) is that a change in one dimension corresponds to a change in one factor of variation in the underlying model of the data, while having little effect on others. The underlying model is rarely available for interrogation, however, which makes learning disentangled representations challenging. Several excellent works for disentangled representation have been proposed that focus on enforcing independence over the latent variables that control the factors of variation (Higgins et al., 2017; Chen et al., 2018; Locatello et al., 2019; Kim and Mnih, 2018; Locatello et al., 2020). In many applications, however, the latent variables are not statistically independent, which is at odds with the notion of disentanglement, i.e., foot length and body height exhibit strong correlation in the observed data (Träuble et al., 2021).
Causal representation learning avoids the aforementioned limitation, as it aims to learn a representation that exposes the unknown high-level causal structural variables, and the relationships between them, from a set of low-level observations (Schölkopf et al., 2021). Unlike disentangled representation learning, it identifies the possible causal relations among latent variables. In fact, disentangled representation learning can be viewed as a special case of causal representation learning where the latent variables have no causal influences (Schölkopf et al., 2021). One of the most prominent additional capabilities of causal representations is the ability to represent interventions and to make predictions regarding such interventions (Pearl, 2000), which enables the generation of new samples that do not lie within the distribution of the observed data. This can be particularly useful to improve the generalization of the resulting model. Causal representations also enable answering counterfactual questions, e.g., would a given patient have suffered heart failure if they had started exercising a year earlier?
Despite its advantages, causal representation learning is a notoriously hard problem—without certain assumptions, identifying the true latent causal model from observed data is generally not possible. There are four primary approaches to achieve identifiability: 1) adapting (weakly) supervised methods with given latent causal graphs or/and labels (Kocaoglu et al., 2018; Yang et al., 2021; Von Kügelgen et al., 2021; Brehmer et al., 2022), 2) imposing special graphical conditions, e.g., with bottleneck graphical conditions (Adams et al., 2021; Xie et al., 2020a; Lachapelle et al., 2021), 3) using temporal information (Yao et al., 2021; Lippe et al., 2022), 4) using hard interventional data (Ahuja et al., 2023; Seigal et al., 2022). A brief review is provided in Section 2. For the supervised approach, when labels are known, the challenging identifiability problem in latent space has been transferred to an identifiability problem in the observed space, for which some commonly-used functional classes have been proven to be identifiable (Zhang and Hyvarinen, 2012; Peters et al., 2014). Given latent causal graphs overly depends on domain knowledge. For the second approach, many true latent causal graphs do not satisfy the assumed special graph structure. The temporal approach is only applicable when temporal information or temporal intervened information among latent factors is available. For using intervention data, most studies focus on hard interventions, particularly single-node and paired hard interventions. In contrast, our work considers the change of causal influences, which is more aligned with soft interventions.
In this work, we explore a new direction in the identifiability of latent causal representations by examining the change in causal influences among latent causal variables, motivated by recent advances in nonlinear ICA (Hyvarinen et al., 2019; Khemakhem et al., 2020). Hyvarinen et al. (2019); Khemakhem et al. (2020) have shown that with an additional observed variable $\mathbf{u}$ to modulate latent independent variables, the latent independent variables conditional on $\mathbf{u}$ are identifiable. This raises a natural question: beyond the independent latent variables in nonlinear ICA, with causal relationships among latent variables, what additional assumptions are required for identifiability of latent causal variables? To answer the question above, we start from the analysis of three intrinsic indeterminacies in latent space (see Section 3): transitivity, permutation indeterminacy, and scaling indeterminacy, which further give rise to the following insights. 1) Transitivity is the scourge of identifiability of latent causal models. 2) Permutation indeterminacy nature enables us to pre-define a causal order, avoiding troublesome directed acyclic graph (DAG) constraints. 3) Scaling and permutation indeterminacy only allow recovering latent causal variables up to permutation and scaling, not the exact values. To overcome the challenge of transitivity, we model the underlying causal representation with weight-variate linear Gaussian models (see Section 4). In this model, both the weights (i.e., causal influences) and the mean and variance of the Gaussian noise are modulated by an additional observed variable $\mathbf{u}$. By exploring the change in causal influences, we demonstrate that the latent causal representations can be recovered up to a trivial permutation and scaling. Combined with the independent causal mechanism, a common assumption in causality, we show that the latent causal structure is also identifiable. Further, realizing that requiring all weights to change might be challenging in real applications, we show partial identifiability results if only part of the weights change in Section 5. Based on our analysis, in Section 6, we additionaly propose a novel method, Structural caUsAl Variational autoEncoder (SuaVE), for learning latent causal representations and latent causal graph among causal representations. Section 7 verifies the efficacy of the proposed approach on both synthetic and real fMRI data.
The key to the identifiability result in this work lies in exploring the changes in causal influences (weights) among the latent causal variables. Essentially, these changes correspond to distribution shifts, which can be viewed as natural interventions acting on latent causal variables in the context of causal representation learning. Such shifts frequently occur in various fields, including medical imaging (Chandrasekaran et al., 2021), biogeography (Pinsky et al., 2020), and finance (Gibbs and Candes, 2021). Similar to the exploration of distribution shifts in observed space (Ghassami et al., 2018; Huang et al., 2020), distribution shifts in latent space also provide opportunities for comparative analysis, offering insights into the underlying causal mechanisms governing the relationships between latent variables. Therefore, investigating distribution shifts for the identifiability of causal representations is a promising approach.
2 Related work
Due to the challenges of identifiability in causal representation learning, most existing works handle this problem by imposing assumptions. We thus give a brief review of the related work on this basis.
(Weakly) Supervised Causal Representation Learning
Approaches falling within this class assume known latent causal graphs or labels. CausalGAN (Kocaoglu et al., 2018) requires a-priori knowledge of the structure of the causal graph of latent variables, which is a significant practical limitation. CausalVAE (Yang et al., 2021) needs additional labels to supervise the learning of latent variables. Such labels are not commonly available, however, and manual labeling can be costly and error-prone. Von Kügelgen et al. (2021) use a known but non-trivial causal graph between content and style factors to study self-supervised causal representation learning. Brehmer et al. (2022) learn causal representation in a weakly supervised setting whereby they assume access to data pairs representing the system before and after a randomly chosen unknown intervention.
Sparse Graphical Structure
Most recent progress in identifiability focuses on sparse graphical structure constraints (Silva et al., 2006; Shimizu et al., 2009; Anandkumar et al., 2013; Frot et al., 2019; Cai et al., 2019; Xie et al., 2020a, 2022a). Adams et al. (2021) provided a unifying viewpoint of this approach whereby a sparser model that fits the observation is preferred. However, they primarily consider linear relations between latent causal variables, and between latent causal variables and observed variables, which are often violated in real-world applications. Lachapelle et al. (2021) handles nonlinear causal relation between latent variables by assuming special sparse graphical structures. However, many latent causal graphs in reality may be more or less arbitrary, beyond a purely sparse graph structure. In contrast, our work assumes a function class of latent variables, and does not restrict the graph structure over them.
Temporal Information
The temporal constraint that the effect cannot precede the cause has been used repeatedly in latent causal representation learning (Yao et al., 2021; Lippe et al., 2022; Yao et al., 2022). For example, Yao et al. (2021) recover latent causal variables and the relations between them using Variational AutoEncoders and enforcing constraints in causal process prior. Lippe et al. (2022) learn causal representations from temporal sequences, which requires the underlying causal factors to be intervened. All of these works can be regarded as special cases of exploring the change of causal influences among latent variables in time series data. The approach proposed is more general since the observed auxiliary variable $\mathbf{u}$ could represent time indices, domain indices, or almost any additional or side information.
Hard Intervention
Very recently, there have been some works exploring interventional data (Brehmer et al., 2022; Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023). Most of them consider hard intervention, and more restricted single-node and paired hard interventions (Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023, 2024; Buchholz et al., 2024; von Kügelgen et al., 2024), which can only capture some special changes of causal influences. By contrast, this work studies unpaired data, and employs soft interventions to model a broader range of possible changes. This is especially important considering that distribution shifts in real applications may be arbitrary. From this perspective, soft interventions could be easier to achieve for latent variables than hard interventions.
Besides, Kivva et al. (2021) considers a nonlinear setting for latent causal discovery. It assumes discrete latent causal variables that are rendered identifiable by a mixture oracle, while here we consider continuous latent causal variables.
3 Indeterminacies in Latent Causal Models
In this section, we first build a connection between nonlinear ICA and causal representation learning, by exploiting the correspondence between the independence of latent variables in nonlinear ICA and the independence of latent noise variables in causal representations. We then consider three indeterminacies in latent space: transitivity, permutation indeterminacy, and scaling indeterminacy, and analyze their impact on identifiability.
3.1 Relating Causal Representation with Identifiable Nonlinear ICA
Figure 1: Causal representation with $\mathbf{u}$.
Causal representation learning aims to uncover latent higher-level causal representations that can explain lower-level raw observations with a nonlinear mapping. Specifically, we assume that the observed variables $\mathbf{x}$ are influenced by latent causal variables ${z_{i}}$, and the causal structure among ${z_{i}}$ can be any directed acyclic graph (which is unknown). For each latent causal variable $z_{i}$, there is a corresponding latent noise variable $n_{i}$, as shown in Figure 1, which represents some unmeasured factors that influence $z_{i}$. The latent noise variables $n_{i}$ are assumed to be independent with each other, conditional on the observed variable $\mathbf{u}$ [^1] For convenience in the later parts, with a slight abuse of definition, independent $n_{i}$ means that $n_{i}$ are mutually independent conditional on the observed variable $\mathbf{u}$., in a causal system (Peters et al., 2017), so it is natural to leverage recent progress in nonlinear ICA (Hyvarinen et al., 2019; Khemakhem et al., 2020), which has shown that the independent latent noise variables $n_{i}$ are identifiable under relatively mild assumptions, i.e., one of the main assumptions is that $\{n_{i}\}$ are Gaussian distributed with their mean and variance modulated by an observed variable $\mathbf{u}$. Taking one step further, our goal is to recover the latent causal variables $z_{i}$. However, with the assumptions for the identifiability of latent noise variables $n_{i}$ from nonlinear ICA, it is still insufficient to identify the latent causal variables $z_{i}$. The reason of such non-identifiability will be given in Section 3.2. Given this fact, a further question is what additional conditions are needed to recover the latent causal variables $z_{i}$. The corresponding identifiability conditions will be given in Section 4.
Understanding $\mathbf{u}$
Essentially, the introduced $\mathbf{u}$ as a surrogate variable is to characterize the changes of independent noise variables $n_{i}$ conditional on $\mathbf{u}$. In some cases, such a surrogate variable may be unavailable or hard to obtain in practice Liu et al. (2021); Creager et al. (2021). Fortunately, it has been shown that learning or inferring such a surrogate variable is possible with additional auxiliary information Lin et al. (2022). These information is often cheaply available for every input in practice Xie et al. (2020b); Wang et al. (2020). Examples include time index of the data in time series forecasting tasks Mudelsee (2019), locations (longitude and latitude) of collected satellite data in remote sensing Rußwurm et al. (2020), modality index in multi-modality dataset, and label in natural images.
3.2 Transitivity: the Challenge of Identifying Causal Representations
Even with the identifiable $n_{i}$, it is impossible to identify the latent causal variables $z_{i}$ without additional assumptions. To interpret this point, for simplicity, let us only consider the influences of $z_{1}$ and $z_{2}$ on $\mathbf{x}$ in Figure 1. According to the graph structure in Figure 1, assume that $z_{1}:=n_{1}$, $z_{2}:=z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})+{\boldsymbol{\varepsilon}}$ (case 1). We then consider the graph structure shown in the left column of Figure 2, where the edge $z_{1}\rightarrow z_{2}$ has been removed, and assume that $z_{1}:=n_{1}$, $z_{2}:=n_{2}$ and $\mathbf{x}:=\mathbf{f}\circ\mathbf{g}(z_{1},z_{2})+{\boldsymbol{\varepsilon}}$ where $\mathbf{g}(z_{1},z_{2})=[z_{1},z_{1}+z_{2}]$ (case 2). Interestingly, we find that the causal models in case 1 and case 2 generate the same observed data $\mathbf{x}$, which implies that there are two different causal models to interpret the same observed data. Clearly, $z_{2}$ in both the two equivalent causal structures are different, and thus $z_{2}$ is unidentifiable.
Figure 2: Two equivalent structures.
Similarly, we can cut the edge $z_{1}\rightarrow z_{3}$ in Figure 1, obtaining another equivalent causal graph as shown in the right of Figure 2. That is, we can have two different $z_{3}$ to interpret the same observed data, and thus $z_{3}$ is unidentifiable. Since there exist many different equivalent causal structures, the latent causal variables in Figure 1 are unidentifiable. Such a result is because the effect of $z_{1}$ on $z_{2}$ (or $z_{3}$) in latent space can be ‘absorbed’ by the nonlinear function from $\mathbf{z}$ to $\mathbf{x}$. We term this indeterminacy transitivity in this work. We will show how to handle this challenge in Section 4. Before that, we introduce the other two indeterminacies in latent space, which assists in understanding our identifiability result.
3.3 Scaling Indeterminacy in Causal Representations
The scaling indeterminacy of latent causal variables is also an intrinsic indeterminacy in latent space. Again, for simplicity, we only consider the influence of $z_{1}$ and $z_{2}$ on $\mathbf{x}$ in Figure 1, and assume that $z_{1}:=n_{1}$, $z_{2}:=z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})$. Under this setting, if the value of $z_{1}$ is scaled by $s$, e.g. $s\times z_{1}$, we can easily obtain the same observed data $\mathbf{x}$ by: 1) letting $z_{2}:=\frac{1}{s}\times z_{1}+n_{2}$ and 2) $\mathbf{x}:=\mathbf{f}\circ\mathbf{g}(z_{1},z_{2})$ where $\mathbf{g}(z_{1},z_{2})=[s\times z_{1},z_{1}]$. This indeterminacy is because the scaling of the latent variables $z_{i}$ can be ‘absorbed’ by the nonlinear function from $\mathbf{z}$ to $\mathbf{x}$ and the causal functions among the latent causal variables. Therefore, without additional information to determine the values of the latent causal variables $z_{i}$, it is only possible to identify the latent causal variable up to scaling, not exactly recovering the values. In general, this scaling does not affect identifying the causal structure among the latent causal variables. We will further discuss this point in Section 4.
3.4 Permutation Indeterminacy in Causal Representations
Due to the nature of ill-posedness, latent causal representation learning suffers from permutation indeterminacy, where the recovered latent causal representations have an arbitrary permutation. For example, assume that the underlying (synthetic) latent causal representations are corresponding to the size, color, shape, and location of an object, and we obtain the recovered latent causal representations $z_{1},z_{2},z_{3},z_{4}$. Permutation indeterminacy means that we can not ensure that the recovered $z_{1}$ represents which specified semantic information, e.g., the size or the color. Therefore, without additional information, it is only possible to identify the latent causal representations up to permutation.
Figure 3: A Causal Fully-connected Graph, based on pre-defined causal order $z_{1},z_{2},z_{3},z_{4}$, thanks to permutation indeterminacy in latent space.
However, the flexibility and ambiguity above in latent space also offer an advantage, i.e., by pre-defining a causal order, we can avoid constraints like those proposed in Zheng et al. (2018), facilitating the learning of latent causal representations and causal structure. For example, For example, assume that the true latent causal relations for generating data are: size $\rightarrow$ color $\rightarrow$ shape $\rightarrow$ location. During the learning process, we can pre-define a causal order $z_{1},z_{2},z_{3},z_{4}$, without specifying semantic information for the nodes. Since $z_{1}$ is the first node in the predefined causal order, $z_{1}$ is enforced to the semantic information of the first node in the true underlying causal order, e.g., the size. Similarly, $z_{2}$ is enforced to the semantic information of the second node in the true underlying causal order, e.g., the color. Furthermore, the predefined causal order allows us to naturally establish a causal fully-connected graph as depicted in Figure 3, ensuring the estimation of a directed acyclic graph (DAG) in learning causal representation and avoiding DAG constraints, such as those proposed by Zheng et al. (2018). We will further demonstrate how to implement this in the proposed method in Section 6.
4 Identifiable Causal Representations with Weight-variant Linear Gaussian Models
As discussed in Section 3.2, the key factor that impedes identifiable causal representations is the transitivity in latent space. Note that the transitivity is because the causal influences among the latent causal variables may be ‘absorbed’ by the nonlinear mapping from latent variables $\mathbf{z}$ to the observed variable $\mathbf{x}$. To address this issue, motivated by identifying causal structures with the change of causal influences in observed space (Ghassami et al., 2018; Huang et al., 2019, 2020), we allow causal influences among latent causal variables to be modulated by an additionally observed variable $\mathbf{u}$, represented by the red edge in Figure 4.
Figure 4: The proposed Latent Causal Model. Here we use the ’red’ lines to indicate the changes of weights across $\mathbf{u}$, which is different from the definition of edges (e.g., causal direction) in standard SCM.
Intuitively, variant causal influences among latent causal variables cannot be ‘absorbed’ by an invariant nonlinear mapping from $\mathbf{z}$ to $\mathbf{x}$, breaking the transitivity, resulting in identifiable causal representations. Specifically, we explore latent causal generative models where the observed data $\mathbf{x}$ is generated by the latent causal variables $\mathbf{z}$, allowing for any potential graph structures among $\mathbf{z}$. In addition, there exist latent noise variables $\mathbf{n}$, known as exogenous variables in causal systems, corresponding to latent causal variables. We introduce a surrogate variable $\mathbf{u}$ characterizing the changes in the distribution of $\mathbf{n}$, as well as the causal influences among latent causal variables $\mathbf{z}$. Here $\mathbf{u}$ could be environment, domain, or time index. More specifically, we parameterize the causal generative models as follows:
$n_{i}:\sim{\cal N}({\beta_{i,1}(\mathbf{u})},{\beta_{i,2}(\mathbf{u})}),$ (1)
${z_{i}}:={\boldsymbol{\lambda}^{T}_{i}(\mathbf{u})(\mathbf{z})}+n_{i},$ (2)
$\mathbf{x}:=\mathbf{f}(\mathbf{z},\boldsymbol{\varepsilon}).$ (3)
where
  • each noise term $n_{i}$ is Gaussian distributed with mean ${\beta_{i,1}(\mathbf{u})}$ and variance ${\beta_{i,2}(\mathbf{u})}$, $\mathbf{n}\in\mathbb{R}^{n}$; both ${\beta_{i,1}}$ and ${\beta_{i,2}}$ can be nonlinear mappings. Moreover, the distribution of $n_{i}$ is modulated by the observed variable $\mathbf{u}$.
  • In Eq. (2), $\boldsymbol{\lambda_{i}}(\mathbf{u})$ denote the vector corresponding to the causal weights from $\mathbf{z}$, $\mathbf{z}\in\mathbb{R}^{n}$, to $z_{i}$, e.g. $\boldsymbol{\lambda_{i}}(\mathbf{u})=[\lambda_{1,i}(\mathbf{u}),\lambda_{2,i}(\mathbf{u}),...,\lambda_{n,i}(\mathbf{u})]$, and each ${\lambda_{j,i}}$ could be a nonlinear mapping.
  • In Eq. (3), $\mathbf{f}$ denote a nonlinear mapping from $\mathbf{z}$ to $\mathbf{x}$, $\mathbf{x}\in\mathbb{R}^{d}$ and $\boldsymbol{\varepsilon}$ is independent noise with probability density function $p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})$, $\boldsymbol{\varepsilon}\in\mathbb{R}^{d-n}$.
In Eq. (1), we assume the latent noise variables to be Gaussian, thus the joint distribution can be expressed by the following exponential family:
$p_{\mathbf{(\mathbf{T}_{\mathbf{n}},\boldsymbol{\beta})}}({\mathbf{n}|\mathbf{u}})=\frac{1}{Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u})}\exp{{\Big{(}\mathbf{T_{n}}^{T}(\mathbf{n})\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u})\Big{)}}},$ (4)
where $Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u})$ denotes the normalizing constant, and $\mathbf{T}_{\mathbf{n}}(\mathbf{n})$ denotes the sufficient statistic for $\mathbf{n}$, whose the natural parameter $\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u})$ depends on $\beta_{i,1}$ and $\beta_{i,2}$.
In Eq. (2), we assume the latent causal model of each $z_{i}$ satisfies a linear causal model with the causal weights being modulated by $\mathbf{u}$, and we call it weight-variant linear Gaussian model. Therefore, $p({\mathbf{z}|\mathbf{u}})$ satisfies the following multivariate Gaussian distribution:
$p_{\mathbf{(\boldsymbol{\lambda},\boldsymbol{\beta})}}({\mathbf{z}|\mathbf{u}})={\cal N}(\boldsymbol{\mu},\Sigma),$ (5)
with the mean $\boldsymbol{\mu}$ and the covariance matrix $\Sigma$ computed by the following recursion relations (Bishop and Nasrabadi, 2006; Koller and Friedman, 2009):
${\mu}_{i}=\sum_{j\in\mathrm{pa}_{i}}\lambda_{j,i}(\mathbf{u}){\mu}_{j}+{\beta_{i,1}(\mathbf{u})},$ (6)
$\Sigma_{i,i}=\sum_{j\in\mathrm{pa}_{i}}\lambda^{2}_{j,i}(\mathbf{u})\Sigma_{j,j}+{\beta_{i,2}(\mathbf{u})},$
$\Sigma_{i,j}=\sum_{k\in\mathrm{pa}_{j}}\lambda_{k,j}(\mathbf{u})\Sigma_{i,k},\quad for\quad i\neq j,$
where $\mathrm{pa}_{i}$ denotes the parent nodes of $z_{i}$. Here, $\mathbf{u}$ can represent the domain index, condition index, or time index. Different values of $\mathbf{u}$ give different causal models of $\mathbf{z}$, resulting in nonstationary data of $\mathbf{z}$, so the modulation of $\mathbf{u}$ can be considered as soft interventions done by nature. Making use of distribution shifts to identify causal structures has been widely considered for causal structure learning among observed variables (Ghassami et al., 2018; Huang et al., 2019, 2020; Schölkopf et al., 2021), and here we extend it to identify latent causal structures.
Furthermore, this multi-variate Gaussian can be re-formulated with the following exponential family:
$p_{\mathbf{(\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta})}}({\mathbf{z}|\mathbf{u}})=\frac{1}{Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u})}\exp{{\Big{(}\mathbf{T}^{T}_{\mathbf{z}}(\mathbf{z})\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u})\Big{)}}},$ (7)
where the parameter $\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u})=[{\Sigma}^{-1}\boldsymbol{\mu};-\frac{1}{2}\operatorname{vec}({{\Sigma}^{-1}})]$, $Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u})$ denotes the normalizing constant, and the sufficient statistic $\mathbf{T}_{\mathbf{z}}(\mathbf{z})=[\mathbf{z};\operatorname{vec}(\mathbf{z}\mathbf{z}^{T})]$ (‘vec’ denotes the vectorization of a matrix). We further denote by $\mathbf{T}_{\text{min}}$ the minimal sufficient statistic and by $k$ its dimension, with $22n\leq k\leq n+(n(n+1))/$ 2. In particular, $k=2n$ corresponds to the case when $\mathbf{T}_{\text{min}}(\mathbf{z})=[{z_{1}},z_{2},...,z_{n},{z^{2}_{1}},{z^{2}_{2}},...,{z^{2}_{n}}]$, or in other words, there is no edges among the latent variables $\mathbf{z}$, while $k=n+(n(n+1))/2$ corresponds to a full-connected causal graph over $\mathbf{z}$. So with different causal structures over $\mathbf{z}$, the dimension $k$ may vary. For the graph structure in Figure 1, $\mathbf{T}_{\text{min}}(\mathbf{z})=[{z_{1}},z_{2},z_{3},z_{4},{z^{2}_{1}},{z^{2}_{2}},{z^{2}_{3}},{z^{2}_{4}},z_{1}z_{2},z_{1}z_{3},z_{3}z_{4}]$ and $k=11$.
The following theorem shows that under certain assumptions on the nonlinear mapping $\mathbf{f}$ and the variability of $\mathbf{u}$, the latent variable $\mathbf{z}$ can be identifiable up to trivial permutation and scaling.
Theorem 1
Suppose latent causal variables $\mathbf{z}$ and the observed variable $\mathbf{x}$ follow the generative models defined in Eq. (1)- Eq. (3), with parameters $({\mathbf{f},\boldsymbol{\lambda},\boldsymbol{\beta}})$. Assume the following holds:
  • (i)
    $\boldsymbol{\varepsilon}$ is independent noise with probability density function $p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})$, which is always finite.
  • (ii)
    The function $\mathbf{f}$ in Eq. (3) is invertible and smooth.
  • (iii)
    There exist $12n+$ 1 distinct points $\mathbf{u}_{\mathbf{n},0},\mathbf{u}_{\mathbf{n},1},...,\mathbf{u}_{\mathbf{n},2n}$ such that the matrix
    $\mathbf{L}_{\mathbf{n}}=(\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},1})-\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},0}),...,\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},2n})-\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},0}))$ (8)
    of size $2n\times 2n$ is invertible.
  • (iv)
    There exist $k+1$ distinct points $\mathbf{u}_{\mathbf{z},0},\mathbf{u}_{\mathbf{z},1},...,\mathbf{u}_{\mathbf{z},k}$ such that the matrix
    $\mathbf{L}_{\mathbf{z}}=(\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},1})-\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},0}),...,\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},k})-\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},0}))$ (9)
    of size $k\times k$ is invertible.
  • (v)
    The function class of $\lambda_{i,j}$ can be expressed by a Taylor series: for each $\lambda_{i,j}$, $\lambda_{i,j}(\mathbf{u}=\mathbf{0})=0$,
then the true latent causal variables $\mathbf{z}$, which are learned by matching the true marginal data distribution $p(\mathbf{x}|\mathbf{u})$, are related to the estimated latent causal variables ${\mathbf{\hat{z}}}$ by the following relationship: $\mathbf{z}=\mathbf{P}\mathbf{\hat{z}}+\mathbf{c},$ where $\mathbf{P}$ denotes the permutation matrix with scaling, $\mathbf{c}$ denotes a constant vector.
Proof sketch
The proof can be done according to the following intuition. With the support of assumptions (i)-(iii), one can utilize the identifiability result from nonlinear ICA to identify the latent noise variables $\mathbf{n}$ up to permutation and scaling, e.g., $\mathbf{n}=\mathbf{P}\mathbf{\hat{n}}+\mathbf{c}_{\mathbf{n}}$ where $\mathbf{\hat{n}}$ denotes the recovered latent noise variables obtained by matching the true marginal data distribution, and $\mathbf{c}_{\mathbf{n}}$ notes a constant vector. Next, using a similar technique, we can utilize assumption (iv) to establish a linear transformation between the true latent causal variables $\mathbf{z}$ and the recovered latent causal variables $\mathbf{\hat{z}}$, e.g., $\mathbf{z}=\mathbf{A}(\mathbf{\hat{z}})+\mathbf{c}_{\mathbf{z}}$ where $\mathbf{c}_{\mathbf{z}}$ denotes a constant vector, and $\mathbf{A}$ denotes a liner transformation. Finally, combining the identifiability result for the latent noise variables $\mathbf{n}$ with assumption (v), the linear transformation can be further reduced to permutation and scaling, e.g., $\mathbf{z}=\mathbf{P}(\mathbf{\hat{z}})+\mathbf{c}_{\mathbf{z}}$. Details can be found in Appendix A.1.
Assumptions (i)-(iv)
Assumptions (i)-(iii) are originally developed by nonlinear ICA Hyvarinen and Morioka (2016); Hyvarinen et al. (2019); Khemakhem et al. (2020); Sorrenson et al. (2020). We here consider unitizes these assumptions considering the following two main reasons. 1) These assumptions have been verified to be practicable in diverse real-world application scenarios Kong et al. (2022); Xie et al. (2022b). 2) Our result eliminates the need to make assumption about the dimensionality of latent causal or noise variables, which is in contrast to existing methods that require prior knowledge of the dimensionality, due to imposing the two-parameter Gaussian noise variables Sorrenson et al. (2020). Assumptions (iv)-(v) originally introduced by this work. Intuitively, assumption (iv) is similar to assumption (iii) in the hope that the auxiliary variable $\mathbf{u}$ must have a sufficiently strong and diverse effect on the distributions of latent causal variables $\mathbf{z}$, similar to latent noise variables $\mathbf{n}$.
Assumptions (v) and sufficient change
Among the assumptions above, the only one that cannot be considered weak or natural is assumption (v), which was originally introduced in this work. Essentially, it is to ensure sufficient changes in causal influence. For simplicity, consider two variables, $z_{1}$ and $z_{2}$, and $z_{2}:=\lambda(\mathbf{u})z_{1}+n_{2}$, where $\lambda(\mathbf{u})=\hat{\lambda}(\mathbf{u})+b$, with $\lambda$ and $\hat{\lambda}$ belonging to the same function class. As a consequence, while the causal influence $\lambda(\mathbf{u})$ changes as a whole across $\mathbf{u}$, there always exists a part $bz_{1}$ that remains unchanged across $\mathbf{u}$. This unchanged term $bz_{1}$ across $\mathbf{u}$ can be absorbed into $\mathbf{f}$ the mapping from $\mathbf{z}$ to $\mathbf{x}$, resulting in a possible solution $\hat{z}_{2}:=\hat{\lambda}(\mathbf{u})z_{1}+n_{2}$, instead of the groundtruth $(\hat{\lambda}(\mathbf{u})+b)z_{1}+n_{2}$, which leads to an unidentifiable outcome. Essentially, this unidentifiable result occurs because a part of a weight remains unchanged across $\mathbf{u}$, a situation we refer to as insufficient change. The objective of assumption (v) is to prevent such insufficient changes. Specifically, by constraining the function class of $\lambda$, e.g., $\lambda(\mathbf{u}=0)=0$ and thus $\hat{\lambda}(\mathbf{u}=0)=0$, $\lambda(\mathbf{u})$ can not be expressed by $\hat{\lambda}(\mathbf{u})+b$ with $b\neq 0$. Intuitively, assumption (v) suggests that there is a scenario where an existing edge can be removed at a certain $\mathbf{u}$, e.g., $\lambda(\mathbf{u}=\mathbf{0})=0$. By removing the edge, we can ensure that no part of the weight corresponding to the edge remains unchanged. It is important to emphasize that assumption (v) is to limit the function class, so that $\lambda$ do not include a constant term, such as $b$, by assuming $\lambda(\mathbf{u}=0)=0$. Once samples are drawn from this limited function class, the assumption is met, making the observed data generated by these samples effective for inferring latent causal variables. Therefore, we do not need the samples to necessarily include the specific point $\mathbf{u}=0$, to generate the corresponding observed data, which is then used for inferring the latent causal variables.
Assumption (v) as a Bridge Between Soft and Hard Interventions for Identifiability
Interestingly, assumption (v) suggests that even if an edge exists between two latent causal variables, there can be a scenario where this edge is effectively removed for a specific $\mathbf{u}$, e.g., when $\lambda(\mathbf{u}=\mathbf{0})=0$. This concept closely aligns with the idea of hard interventions, where manipulating a variable can sever causal connections. This insight draws a compelling link between this work and recent advances that leverage hard interventions to improve the identifiability of causal representations (Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023). However, unlike these works, which typically rely on setting $\mathbf{u}=\mathbf{0}$, this work, through assumption (v), defines a restricted class of functions for $\lambda(\mathbf{u})$. By sampling from this restricted function class, identifiability can be achieved without necessarily requiring $\mathbf{u}=\mathbf{0}$, thereby broadening the applicability beyond existing hard intervention methods. Further exploration of this connection could pave the way for a more general framework that unifies soft and hard interventions for identifiability. However, such an investigation lies beyond the scope of this work.
Potential application
Assumption (v), for instance, could arise in the analysis of imaging data of cells, where various batches of cells are exposed to different small-molecule compounds (e.g., $\mathbf{u}$). In this scenario, each latent variable (e.g., $\mathbf{z}$) represents the concentration level of a distinct group of proteins, with protein-protein interactions (e.g., latent causal relations among $\mathbf{u}$) playing a significant role. There is considerable interest in computational methods aimed at extracting biological information from images (e.g., $\mathbf{x}$), given that imaging data offers a more cost-effective alternative to other assays while still yielding valuable insights into cell states Chandrasekaran et al. (2021). Research has revealed that the mechanisms of action of small molecules exhibit variations in selectivity Forbes and Krueger (2019), which can profoundly affect protein-protein interactions. In the context above, assumption (v) can be justifiable as certain small molecules have been demonstrated to significantly influence protein-protein interactions, thereby altering the dependency of the concentration of the products of these interactions on the concentration of the reactants. The assumption require that there exist $\lambda_{i,j}(\mathbf{u}=\mathbf{0})=0$. This implies that the causal relationship may be disconnected. In the context above, it is possible for a small molecule compound to disrupt or inhibit protein-protein interactions (PPIs), effectively causing the interaction to disappear Arkin and Wells (2004). Such small molecules are often referred to as inhibitors or disruptors of protein-protein interactions. These compounds can work through various mechanisms, such as blocking the binding interface between the two proteins, destabilizing the protein complex, altering the conformation of one or both proteins involved in the interaction, or interfering with the signaling pathways associated with the interaction. The development of small molecule inhibitors of PPIs is an active area of research in drug discovery, especially in diseases where dysregulated protein interactions contribute to pathology Lu et al. (2020); Bojadzic et al. (2021).
Insights
Intuitively, allowing causal influences among latent causal variables to change leads to varying causal influences. Such variations cannot be ’absorbed’ by an invariant nonlinear mapping from $\mathbf{z}$ to $\mathbf{x}$, which breaks the transitivity mentioned in Section 3.2, resulting in identifiability. For example, assume that $z_{1}:=n_{1}$, $z_{2}:=\lambda(\mathbf{u})z_{1}+n_{2}$ and $\mathbf{x}:=\mathbf{f}(z_{1},z_{2})$. If we allow the causal influence $\lambda(\mathbf{u})z_{1}$ to be ’absorbed’ by the mapping $\mathbf{f}$, we would have $z_{1}:=n_{1}$, $z_{2}:=n_{2}$, and the resulting mapping, $\mathbf{f^{\prime}}=\mathbf{f}\circ\mathbf{g}$ where $\mathbf{g}(z_{1},z_{2})=[z_{1},\lambda(\mathbf{u})z_{1}+z_{2}]$, would need to include the original $\mathbf{f}$ as well as the term $\lambda(\mathbf{u})$ in $\mathbf{g}$, to obtain the same obervation $\mathbf{x}$. This means that the resulting mapping $\mathbf{f^{\prime}}$ would change across $\mathbf{u}$ due to the term $\lambda(\mathbf{u})$, which violates the fact that the mapping from $\mathbf{z}$ to $\mathbf{x}$ remains unchanged across $\lambda(\mathbf{u})$ as defined in Eq. (3).
Identifiability of Causal Structures among Latent Variables
Theorem 1 has shown that latent causal variables can be identified up to trivial permutation and linear scaling. With this result, the identifiability of causal structure in the latent space reduces to the identifiability of the causal structure in the observational space. Denote by $G_{\mathbf{z}}^{u}$ the causal graph over latent variables $\mathbf{z}$ when $\mathbf{u}=u$, and let $G_{\mathbf{z}}^{\text{union}}=\cup_{u}G_{\mathbf{z}}^{u}$ be the union of causal graphs across different values of $\mathbf{u}$. Similarly, we define $G_{\mathbf{z}\cup\mathbf{u}}^{\text{union}}$ to be the union of causal graphs over $\mathbf{z}\cup\mathbf{u}$. With the help of recent progress in causal discovery from heterogeneous data (Huang et al., 2020), the following corollary shows that the causal structure among latent variables $\mathbf{z}$ is identifiable up to the Markov equivalence class of $G_{\mathbf{z}}^{\text{union}}$. Moreover, although there exists scaling indeterminacy for the recovered latent variables as stated in Theorem 1, it does not affect the identifiability of the causal structure.
Corollary 2
Suppose latent causal variables $\mathbf{z}$ and the observation $\mathbf{x}$ follow the generative models defined in Eq. (1)- Eq. (3), and that the conditions in Theorem 1 hold. Then, under the assumption that $G_{\mathbf{z}}^{\text{union}}$ is a DAG and that the joint distribution over $\mathbf{z}\cup\mathbf{u}$ is Markov and faithful to $G_{\mathbf{z}\cup\mathbf{u}}^{\text{union}}$, the causal structure among variables $\mathbf{z}$ can be identified up to the Markov equivalence class of $G_{\mathbf{z}}^{\text{union}}$. That is, $\textit{MEC}(G_{\mathbf{z}}^{\text{union}})$ is identifiable, where $\textit{MEC}(\cdot)$ denotes the Markov equivalence class.
Proof Sketch
The proof can be done by first establishing that Theorem 1 identifies the latent causal variables up to trivial permutation and linear scaling, which reduces the problem to identifying the causal structure in the observed space. Then, by leveraging the results from Huang et al. (2020), which ensure identifiability of the causal structure in the observed space under the Markov condition and faithfulness assumption, we can identify the causal structure up to the Markov equivalence class. Next, we show that the Markov equivalence class over $\mathbf{z}$ remains unchanged after removing the domain variable $\mathbf{u}$ and its edges, preserving the graph’s skeleton and directions. Finally, since linear scaling does not affect conditional independence relationships, the identifiability of the causal structure in the latent space is preserved. As a result, the causal structure among the latent variables $\mathbf{z}$ can be identified up to the Markov equivalence class. Please refer to Appendix A.3 for more details.
In addition, with the help of the independent causal mechanisms (ICM) principle (Ghassami et al., 2018; Huang et al., 2020; Schölkopf et al., 2021), the following corollary shows that the causal structure among latent variables $\mathbf{z}$ is fully identifiable.
Corollary 3
Suppose the conditions in Corollary 2 hold. Denote by $\boldsymbol{\theta}_{i}$ the involved parameters in the causal mechanism of $z_{i}$, with $\boldsymbol{\theta}_{i}=(\beta_{i,1};\beta_{i,2};\boldsymbol{\lambda}_{i})$, and denote by $z_{pa(i)}$ the parents of $z_{i}$ in $G_{\mathbf{z}}^{u}$. If $\boldsymbol{\theta}_{i}$ and $\boldsymbol{\theta}_{pa(i)}$ change independently across different values of $\mathbf{u}$ for any $z_{i}$, then the acyclic causal structure among latent variables $\mathbf{z}$ can be fully identified. That is, $G_{\mathbf{z}}^{\text{union}}$ is fully identifiable.
Proof Sketch
To prove this corollary, we first introduce the ICM principle, which states that in a causally sufficient system, the causal modules and their parameters change independently across different domains. This principle holds in the causal direction, meaning that changes in one module do not affect others. However, this independence generally breaks down in the anti-causal direction. Clearly, The scaling indeterminacy for the recovered latent variables does not affect this principle. Please refer to Appendix A.4 for more details.
5 Change of Part of Weights and Partial Identifiability Result
The aforementioned theoretical result necessitates that all weights undergo changes across $\mathbf{u}$, as constrained by the assumption (v) in Theorem 1. However, in practical applications, this assumption may not hold true. Consequently, two fundamental questions naturally arise: Is this assumption necessary for identifiability in the absence of any supplementary assumptions? Alternatively, can we obtain partial identifiability results if only some of the weights change across $\mathbf{u}$? In fact, when part of weights change, we can still provide partial identifiability results, as outlined below.
Theorem 4
Suppose latent causal variables $\mathbf{z}$ and the observed variable $\mathbf{x}$ follow the generative models defined in Eq. (1)- Eq. (3). Under the condition that the assumptions (i)-(iv) in Theorem 1 are satisfied, for each $z_{i}$,
  • (a)
    if $z_{i}$ is a root node or all weights to $z_{i}$, e.g., $\boldsymbol{\lambda}_{i}$ in Eq. (2), meet assumption (v) in Theorem 1, then the true $z_{i}$ is related to the recovered one $\hat{z}_{j}$, obtained by matching the true marginal data distribution $p(\mathbf{x}|\mathbf{u})$, by the following relationship: $z_{i}=s\hat{z}_{j}+c$, where $s$ denotes scaling, $c$ denotes a constant,
  • (b)
    if there exists an unchanged weight from the parent node $z_{j}$ to $z_{i}$, e.g., in ${\lambda}_{j,i}$, in Eq. (2), then $z_{i}$ is unidentifiable,
  • (c)
    if there exists an weight that can be expressed as $\lambda_{j,i}(\mathbf{u})+b$, where b is a non-zero constant, from the parent node $z_{j}$ to $z_{i}$, then $z_{i}$ is unidentifiable.
Proof sketch
This can be proved by the fact that regardless of the assumption (v) in Theorem 1 , two results hold in Theorem 1, i.e., $\mathbf{z}=\mathbf{A}(\mathbf{\hat{z}})+\mathbf{c}_{\mathbf{z}}$, and $\mathbf{n}=\mathbf{P}\mathbf{\hat{n}}+\mathbf{c}_{\mathbf{n}}$. Then using the change of all weights in $\boldsymbol{\lambda}_{i}$, we can prove (a). To prove (b), we demonstrate that it is always possible to construct an alternative solution $\hat{z}_{i}$ by removing the term ${\lambda}_{j,i}z_{j}$, corresponding to unchanged weight, in the true ${z}_{i}$, which is capable of generating the same observations $\mathbf{x}$. As a result, we can show non-identifiable result. To prove (c), similar to (b), we can construct an alternative solution $\hat{z}_{i}$ by removing the term $bz_{j}$, corresponding to the constant $b$, in true $z_{i}$.
Insights
1) The result (b) of Theorem 4 implies the necessity of changing weights, without additional assumptions. This can be proven by the contrapositive of (b), i.e., if $z_{i}$ is identifiable, then there do not exist unchanged weights from the parent nodes to $z_{i}$. 2) The result (c) of Theorem 4 implies the necessity of eliminating the corner case where $\lambda(\mathbf{u})=\lambda^{\prime}(\mathbf{u})+b$, without additional assumptions. Similarly, this can be proven by the contrapositive of (c), i.e., if $z_{i}$ is identifiable, then there do not exist a weight that can be expressed as $\lambda(\mathbf{u})+b$. Essentially, this implies that sufficient changes in weights are required for identifiability. Informally, if $z_{i}=\boldsymbol{\lambda}_{i}^{T}\mathbf{z}+n_{i}$, then sufficient changes mean that $\boldsymbol{\lambda}_{i}^{T}$ do not include non-zero constant across $\mathbf{u}$. 3) Together with result (a) of Theorem 4, Theorem 4 implies that the entire latent space can theoretically be partitioned into two distinct subspaces: one subspace pertains to invariant latent variables, while the other encompasses variant variables. This partitioning may be valuable for applications that focus on learning invariant latent variables to adapt to varying environments, such as domain adaptation or generalization.
Discussion 1: Partial Causal Structure
Partial identifiability of latent causal variables in Theorem 4 does not necessarily guarantee the unique recovery of the corresponding partial latent causal graph structure. However, a probable result can still be achieved. Specifically, if there are no interactions (edges) between the two latent subspaces in the ground truth graph structure, it becomes possible to recover the latent causal structure within the latent variant space. When interactions do exist, examining how they affect the recovery of the latent causal graph structure is an intriguing area for further exploration. Additionally, it is valuable to investigate how such partial results influence the outcomes of interventions and counterfactual inference.
Discussion 2: Parent nodes do not impact children
Perhaps counterintuitively, the partial identifiability result in Theorem 4 suggests that the identifiability of $z_{i}$ remains possible even when its parent nodes are unidentifiable. This is primarily due to the identifiability of all latent noise variables $\mathbf{n}$, regardless of changes in the weights. In this context, for latent linear causal models, all necessary information to recover latent causal variables is encapsulated within these identifiable latent noise variables. As a result, it is possible to recover $z_{i}$, irrespective of the identifiability of its parent nodes.
6 Learning Causal Representations with Weight-Variant Linear Models
Based on the identifiable results above, in this section, we propose a structural causal variational autoencoder (SuaVE) to learn latent causal representations. We first propose a structural causal model prior relating to the proposed weight-variant linear Gaussian model. We then show how to incorporate the proposed prior into the traditional VAE framework (Kingma and Welling, 2013a), together with a variational posterior.
DAG constraints
To address the complexity of combinatorial search, recent methods have proposed exact characterizations of DAGs that enable tackling the problem through continuous optimization techniques (Zheng et al., 2018; Yu et al., 2019; Zheng et al., 2020; He et al., 2021). However, this approach does not guarantee the absence of cycles at any stage of training, and solutions often require post-processing (Zantedeschi et al., 2023). By contrast, as mentioned in Section 3.4, the permutation indeterminacy in latent space allows us to predefine a causal order without specifying semantic information. With the guarantee of identifiability, the nodes in the predefined order are enforced to learn the corresponding semantic information in the true causal order for generating data, ensuring each node learns its designated role. Given this, we can incorporate a predefined causal order in the following prior and posterior models.
Prior Model
According to the causal model for the latent causal variables, e.g., Eq. (1) and Eq. (2), we construct a linear Gaussian prior with weights and noise depending on $\mathbf{u}$. Since we can pre-define the causal order as mentioned above, let us assume the causal order is $z_{1},z_{2},..z_{n}$. Given this, we can create a fully connected causal graph based on the predefined causal order, where each $z_{i}$ has parent nodes $z_{i}^{\prime}$ for $i^{\prime}<i$. Summarizing the above, we propose the following prior model:
$p({\bf{z}}|\mathbf{u})=p(z_{1}|\mathbf{u})\prod\limits_{i=2}^{n}{p({{z_{i}}}|{\bf z}_{<i},\mathbf{u})}=\prod\limits_{i=1}^{n}{{\cal N}(\mu_{z_{i}},\delta^{2}_{z_{i}})},$ (10)
where $\mu_{z_{i}}=\sum\limits_{i^{\prime}<i}\lambda_{i^{\prime},i}(\mathbf{u})z_{i^{\prime}}+\beta_{i,1}(\mathbf{u}),\delta^{2}_{z_{i}}=\beta_{i,2}(\mathbf{u})$, and the corresponding weight matrix $\boldsymbol{\lambda}=[\boldsymbol{\lambda}_{1}(\mathbf{u}),...,\boldsymbol{\lambda}_{n}(\mathbf{u})]$ can be constrained as a upper triangular matrix with zero-value diagonal elements. Again, the proposed prior naturally ensures a directed acyclic graph estimation, avoiding traditional DAG constraints.
Variational Posterior and Objective
The nature of the proposed prior in Eq. (10) gives rise to the following variational posterior:
$q({\bf{z}}|\mathbf{x,u})=q(z_{1}|\mathbf{x,u})\prod\limits_{i=2}^{n}{q({{z_{i}}}|{\bf z}_{<i},\mathbf{x,u})}=\prod\limits_{i=1}^{n}{{\cal N}(\mu^{\prime}_{z_{i}},\delta^{{}^{\prime}2}_{z_{i}})},$ (11)
where $\mu^{\prime}_{z_{i}}=\sum\limits_{i^{\prime}<i}\lambda^{{}^{\prime}}_{i^{\prime},i}(\mathbf{u})z_{i^{\prime}}+\beta^{{}^{\prime}}_{i,1}(\mathbf{x,u}),\delta^{{}^{\prime}2}_{z_{i}}=\beta^{{}^{\prime}}_{i,2}(\mathbf{x,u}).$ Therefore, we can arrive at the following evidence lower bound (ELBO):
$\mathop{}\mathbb{E}_{q(\mathbf{z}|\mathbf{x,u})}(\log p(\mathbf{x}|\mathbf{z},\mathbf{u}))-{D_{KL}}({q}(\mathbf{z|x,u})||p(\mathbf{z}|\mathbf{u})),$ (12)
where ${D_{\text{KL}}}$ denotes Kullback–Leibler divergence, a measure of how one probability distribution ${q}(\mathbf{z|x,u})$ is different from a second, reference probability distribution $p(\mathbf{z}|\mathbf{u})$. We implement assumption (v) and incorporate it into the evidence lower bound by the following objective:
$\mathop{}\mathbb{E}_{q(\mathbf{z}|\mathbf{x,u})}(\log p(\mathbf{x}|\mathbf{z},\mathbf{u}))-{D_{\text{KL}}}({q}(\mathbf{z|x,u})||p(\mathbf{z}|\mathbf{u}))-\gamma_{1}\sum\limits_{i,j}||\lambda_{i,j}(\mathbf{u}=\mathbf{0})||_{1}-\gamma_{2}\sum\limits_{i,j}||\lambda^{\prime}_{i,j}(\mathbf{u}=\mathbf{0})||_{1}.$ (13)
In our implementation, we use one-hot encoding for the additional observed variable $\mathbf{u}$. To enforce assumption (v), we also create new data points with $\mathbf{u}=\mathbf{0}$, which correspond to an output of 00. This allows us to translate assumption (v) into an implemented constraint, such as applying the L1 norm on $\lambda_{i,j}$, as discussed in the last two terms of Eq. 13, where $\gamma_{1}$ and $\gamma_{2}$ are two hyper-parameters used to balance the constraints and the ELBO. Figure 5 depicts the implementation of the proposed method to learn causal representations.
Figure 5: Implementation Framework to learn linear Gaussian causal representations.
Discussion: Relating to hierarchical VAE
Beyond standard VAEs, recent advanced works aim to explore hierarchical VAEs with the hope that they could improve the evidence lower bound in standard VAE and decrease reconstruction error compared to standard VAEs. More importantly, the stack of latent variables in hierarchical VAEs might learn a feature hierarchy, where different layers capture information at varying levels of detail and abstraction (Kingma et al., 2016; Sønderby et al., 2016; Maaløe et al., 2019; Vahdat and Kautz, 2020). To this end, hierarchical VAE models often employ the following generative model: $z_{1}\rightarrow z_{2}\rightarrow...\rightarrow z_{n}\rightarrow\mathbf{x}$. Interestingly, hierarchical VAE models can be regarded as a special case of the proposed prior model in Equation (10), where the fully-connected graph is reduced to the simpler connections used in the generative model of hierarchical VAEs. Essentially, the proposed SuaVE also aims to learn hierarchical latent variables. Due to permutation indeterminacy, we can model the prior or posterior as hierarchical structures. The difference is that the edges among latent variables of the proposed SuaVE are learned adaptively from data, unlike the simple connections in hierarchical VAEs. Importantly, the proposed SuaVE has rigorous theoretical justification for learning hierarchical latent variables and the structures among them, whereas hierarchical VAEs lack such support without further assumptions. Further exploring the relationship between hierarchical VAEs and latent causal models is both interesting and worthwhile.
7 Experiments
7.1 Synthetic Data
We first conduct experiments on synthetic data, generated by the following process: we divide the latent noise variables into $M$ segments, where each segment corresponds to one conditional variable $\mathbf{u}$ as the segment label. Within each segment, we first sample the mean ${\beta}_{i,1}$ and variance ${\beta}_{i,2}$ from uniform distributions $[-2,2]$ and $[0.01,3]$, respectively. We then sample the weights $\boldsymbol{\lambda}_{i,j}$ from uniform distributions $[0.1,2]$. Then for each segment, we generate the latent causal samples according to the generative model in Eq.3. Finally, we obtain the observed data samples $\mathbf{x}$ by an invertible nonlinear mapping. More details can be found in Appendix.
Comparison
We compare the proposed method with identifiable VAE (iVAE) (Khemakhem et al., 2020), $\beta$-VAE (Higgins et al., 2017), CausalVAE(Yang et al., 2021), and vanilla VAE (Kingma and Welling, 2013b). Among them, iVAE has been proven to be identifiable so that it is able to learn the true independent noise variables with certain assumptions. While $\beta$-VAE has no theoretical support, it has been widely used in various disentanglement tasks. Note that both methods assume that the latent variables are independent, and thus they cannot model the relationships among latent variables. To make a fair comparison in the unsupervised setting, we implement an unsupervised version of CausalVAE, which is not identifiable.
Figure 6: Performances of the proposed SuaVE with different values of hyper-parameters $\gamma_{1}$ and $\gamma_{2}$. Implementation details can be found in Appendix A.6.2.
Performance metric
Since the proposed method can recover the latent causal variables up to trivial permutation and linear scaling, we compute the mean of the Pearson correlation coefficient (MPC) to evaluate the performance of our proposed method. The Pearson correlation coefficient is a measure of linear correlation between the true latent causal variables and the recovered latent causal variables. Note that the Pearson coefficient is suitable for iVAE, since it has been shown that iVAE can also recover the latent noise variables up to linear scaling under the setting where the mean and variance of the latent noise variables are changed by $\mathbf{u}$ (Sorrenson et al., 2020). To remove the permutation effect, following Khemakhem et al. (2020), we first calculate all pairs of correlation and then solve a linear sum assignment problem to obtain the final results. A high correlation coefficient means that we successfully identified the true parameters and recovered the true variables, up to component-wise linear transformations.
Ablation Study
We first conduct experiments on the proposed method, varying the hyper-parameters $\gamma_{1}$ and $\gamma_{2}$. Figure 6 shows the results of performance of the proposed SuaVE with different values of hyper-parameters. The performance remains relatively stable on synthetic data. This suggests that on randomly generated synthetic data, the variation in weights generally satisfies assumption (v), which requires avoiding insufficient change—specifically, the scenario where a portion of the weights remains unchanged across $\mathbf{u}$. This stability occurs because the weights are independently and randomly generated without any inherent patterns, making it unlikely that they would fit any specific nonlinear/linear function with a constant term. Consequently, the proposed method naturally avoids the special case where some weights remain unchanged. Therefore, in the subsequent experiments, we set the values of the hyper-parameters $\gamma_{1}$ and $\gamma_{2}$ to zero.
Figure 7: Performances of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent causal variables on synthetic data with different numbers of segments. Implementation details can be found in Appendix A.6.1.
Figure 8: Performances of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent causal variables on synthetic data with different numbers of the latent causal variables. Implementation details can be found in Appendix A.6.
Figure 9: The SHD obtained by the proposed SuaVE and CausalVAE. Since CausalVAE is not identifiable, it obtains fully connected graphs, the SHD of the recovered CPDAG is the same as one of the recovered skeletons. Implementation details can be found in Appendix A.6.
Analysis of Comparison Results
We compared the performance of the proposed SuaVE to some variants of VAE mentioned above. We used the same network architecture for encoder and decoder parts in all these models. In particular, we add a sub-network to model the conditional prior in both iVAE and the proposed SuaVE. We further assign a linear SCM sub-network to model the relations among latent causal variables in the proposed SuaVE. We trained these 5 models on the dataset described above, with different numbers of segments and different numbers of latent causal variables. For each method, we use 5 different random seeds for data sampling. Figure 7 shows the performance on two latent causal variables with different numbers of segment. The proposed SuaVE obtains the score 0.96 approximately for all the different numbers of segment. In contrast, $\beta$-VAE, VAE and CausalVAE fail to achieve a good estimation of the true latent variables, since they are not identifiable. iVAE obtains unsatisfactory results, since its theory only holds for i.i.d. latent variables. Figure 8 shows the performance in recovering latent causal variables on synthetic data with different numbers of the latent causal variables. Figure 9 shows the structural Hamming distance (SHD), a standard metric for comparing graphs via their adjacency matrices, for the recovered skeletons (the undirected version of the graph with all directed edges replaced by undirected edges), causal graphs, and completed partial directed acyclic graphs (CPDAGs, which capture the set of all DAGs that are Markov equivalent), by the proposed SuaVE and CausalVAE. Since CausalVAE is not identifiable, it obtains fully connected graphs, the SHD of the recovered CPDAG is the same as one of the recovered skeletons.
7.1.1 Experiments on Changes of Part of Weights
In this part, we conduct experiments on changes of part of weighs, to verify the results in Theorem 4. To this end, we consider two cases on the following graph: $z_{1}\rightarrow z_{2}\rightarrow z_{3}\rightarrow z4$. The data details are similar to those of the 4-dimensional case mentioned in Appendix A.6.2, except that $z_{3}$ is a parent node of $z_{4}$. Case 1: the weight on the edge of $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, $z_{3}\rightarrow z_{4}$) remains unchanged (e.g., a constant) across $\mathbf{u}$, while the weights on the remaining edges change across $\mathbf{u}$, Case 2: the weight on the edge of $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, $z_{3}\rightarrow z_{4}$) includes a constant part (e.g., $\lambda_{1,2}(\mathbf{u})+b$, $b$ is a non-zero constant.) across $\mathbf{u}$, while the weights on the remaining edges change across $\mathbf{u}$. Again, the network architecture, optimization, and hyper-parameters are the same as used for experiments of Figure 7 (a). Figures 10 and 11 show the performance of the proposed method in the settings corresponding to Case 1 and Case 2 mentioned above. According to MPC, we observe that unchanged weights in Case 1 or weights involving a constant term in Case 2 lead to non-identifiability. Conversely, changing weights contribute to the identifiability of the corresponding nodes. These empirical results align with the partial identifiability results in Theorem 4.
Figure 10: Case 1: Performances of the proposed method with the change of part of weights. The ground truth of the causal graph is $z_{1}\rightarrow z_{2}\rightarrow z_{3}\rightarrow z_{4}$. From top to bottom: keeping weight on $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, and $z_{3}\rightarrow z_{4}$) unchanged. Those results are consistent with the analysis of partial identifiability results in Theorem 4.
Figure 11: Case 2: Performances of the proposed method with the change of part of weights. The ground truth of the causal graph is $z_{1}\rightarrow z_{2}\rightarrow z_{3}\rightarrow z_{4}$. From top to bottom: weight on $z_{1}\rightarrow z_{2}$ (and accordingly, $z_{2}\rightarrow z_{3}$, and $z_{3}\rightarrow z_{4}$) includes a non-zero constant part and changing part, e.g., $\lambda(\mathbf{u})+b$. Those results are consistent with the analysis of partial identifiability results in Theorem 4.
7.1.2 Experiments on i.i.d. $z_{i}$
Our work is closely related to nonlinear ICA, and our assumptions include the assumptions for identifying the Gaussian noise variables. To verify this point, we conduct experiments on the case where any two latent causal variables $z_{i}$ are independent given $\mathbf{u}$. Data details are the same as mentioned in section A.6.2, but we here enforce $\lambda_{i,j}(\mathbf{u})=0$ for each $i,j$. Again, the network architecture, optimization and hyper-parameters are the same as section A.6.2, except for learning rate (here we set $21r=1e-$ 2). Figure 12 shows the performances of the proposed SuaVE and iVAE. We can see that iVAE is slightly better than the proposed SuaVE, both obtain satisfying results in terms of MPC.
Figure 12: Performances of the proposed SuaVE and iVAE in recovering the independent latent causal variables conditional $\mathbf{u}$.
7.1.3 Performance for violated assumption on distribution
To further understand the assumptions of our identifiability result, we conduct experiments in a setting where the assumption on distribution is violated. Table 1 shows the performance when the distribution assumption is violated. We observe that violating the distribution assumption leads to unsatisfactory performance in terms of MPC. The possible reasons for the results with Uniform (Laplace, Gamma) Noise may be: 1) the proposed method enforces Gaussian prior and posterior distributions to approximate non-Gaussian noise, and 2) the settings with Uniform (Laplace, Gamma) noise may be inherently non-identifiable.
Table 1: Performance for violated assumptions. Uniform (Laplace, Gamma) Noise: the distribution of latent noise variables. Unchanged Weights: weights keep unchanged across $\mathbf{u}$. Violated assumption (v): $\lambda_{i,j}({\mathbf{0}})=c$ where $c$ a non-zero constant. Matching Assumptions: a setting that matches our theoretical assumptions.
Generative process MPC(Mean $\pm$ Std)
Uniform Noise 0.83 $\pm$ 0.09
Laplace Noise 0.83 $\pm$ 0.03
Gamma Noise 0.82 $\pm$ 0.05
Matching Assumptions 0.95 $\pm$ 0.01
7.2 Image Data
We further verify the proposed identifiability results and method on images from the chemistry dataset proposed in (Ke et al., 2021), which is corresponding to simple chemical reactions where the state of an element can cause changes to another variable’s state. The environment consists of a number of objects whose positions are kept fixed and thus, uniquely identifiable, while the colors (states) of the objects affected by this variable (according to the causal graph) can change. We use a weight-variant linear Gaussian model as mentioned in section A.6.2 for a 3-dimensional case, to generate the latent variables, and obtain the corresponding observational images. A visualization of the environment can be found in Figure 13.
Figure 13: Demonstration of the chemistry environment. The colors (states) of the objects change according to the causal graph: the ’diamond’ causes the ‘triangle’, and the ‘triangle’ causes the ’square’.
Figure 14 shows the MPC obtained by different methods. The proposed method performs better than $\beta$-VAE(Higgins et al., 2017) and CausalVAE (Yang et al., 2021). In term of MPC, the proposed method successfully learns causal representations, while the others fails to learn latent causal variables. To further verify whether the learned causal structure by the proposed method is consistent with the true underlying latent structure, we investigate the results of intervening on each learned variable. Figure 15 shows the intervention results on each learned latent variable by the proposed SuaVE. We can see from Figure 15 that:
  • Intervention on $z_{1}$ (the ’diamond’) cause the change of both $z_{2}$ and $z_{3}$ (the colors of ’triangle’ and ’square’), as depicted in the left in Figure 15.
  • Intervention on $z_{2}$ (the ’triangle’) only cause the change of $z_{3}$ (the color of ’square’), in the medium in Figure 15.
  • Intervention on the changes of $z_{3}$ (the ’square’) do not cause the change of $z_{1}$ and $z_{2}$ (the ’diamond’ and the ’triangle’), in the right in Figure 15
All these observations are consistent with the ground truth latent causal structure, e.g., $z_{1}\rightarrow z_{2}\rightarrow z_{3}$, together with results in Figure 14, demonstrating that the proposed method successfully learns latent causal structures. Similar, we inversigate the change in reconstructions while traversing each learned latent variable obtained by $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021). Figure 16 and 17 show traversal results on the learned representations by $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021), respectively. we observed that changing each learned variable results in changes to the colors of all objects. Together with the results in Figure 14, these suggests that $\beta$-VAE (Higgins et al., 2017) and CausalVAE (Yang et al., 2021) tend to learn representations that entangle the true underlying latent causal variables and fail to recover the true underlying latent causal structure.
Ours
$\beta$-VAE
CausalVAE
Figure 14: MPC obtained by different methods on the image dataset. Supported by our identifiability analysis, the proposed method achieves satisfactory MPC, indicating its success in recovering latent causal representations.
Figure 15: Intervention results on the learned latent variables by the proposed method SuaVE. From left to right: intervention on the learned $z_{1},z_{2},z_{3}$, respectively. The vertical axis denotes different samples, The horizontal axis denotes enforcing different values on the learned causal representation. We can see that: 1) (from the left) the changes of $z_{1}$ (the ’diamond’) cause the change of both $z_{2}$ and $z_{3}$ (the ’triangle’ and the ’square’). 2) (from the middle) the changes of $z_{2}$ (the ’triangle’) only cause the change of $z_{3}$ (the ’square’). 3) (from the right) the changes of $z_{3}$ (the ’square’) do not cause the change of $z_{1}$ and $z_{2}$ (the ’diamond’ and the ’triangle’).
Figure 16: Traversal results on the learned latent variables by $\beta$-VAE (Higgins et al., 2017). From left to right: Traversal on the learned $z_{1},z_{2},z_{3}$, respectively. The vertical axis denotes different samples, The horizontal axis denotes enforcing different values on the learned causal representation. We can see that from the left to the right with changes of anyone $z_{i}$, the remaining two variables always changes, which means that each learned variable by $\beta$-VAE (Higgins et al., 2017) tend to be mixtures of the true underlying latent causal variables.
Figure 17: Traversal results on the learned latent variables by CausalVAE (Yang et al., 2021). From left to right: Traversal on the learned $z_{1},z_{2},z_{3}$, respectively. The vertical axis denotes different samples, The horizontal axis denotes enforcing different values on the learned causal representation. Again, we can see that from the left to the right with changes of anyone $z_{i}$, the remaining two latent variables also change. These results mean that each learned variable tend to be mixtures of the true underlying latent causal variables, and CausalVAE (Yang et al., 2021) fails to recovering latent causal structure.
7.3 fMRI Data
Following Ghassami et al. (2018), we applied the proposed method to fMRI hippocampus dataset (Laumann and Poldrack, 2015), which contains signals from six separate brain regions: perirhinal cortex (PRC), parahippocampal cortex (PHC), entorhinal cortex (ERC), subiculum (Sub), CA1, and CA3/Dentate Gyrus (DG) in the resting states on the same person in 84 successive days. Each day is considered as different $\mathbf{u}$, thus $\mathbf{u}$ is a 84-dimensional vector. Since we are interested in recovering latent causal variables, we treat the six signals as latent causal variables by applying a random nonlinear mapping on them to obtain observed data. We then apply various methods on the observed data to recover the six signals. Figure 19 (a) shows the performance of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent six signals. $\beta$-VAE aims to recover independent latent variables, and it obtains an interesting result: enforcing independence (e.g., $\beta=25,50$) leads to worse MPC, and relaxing it (though contracting to its own independence assumption) improves the result (e.g. $\beta=4$). This is because the latent variables given the time index are not independent in this dataset.
We further verify the ability of the proposed method to recover causal skeletons. We use the anatomical connections (Bird and Burgess, 2008) as a reference as shown in Figure 19 (b), since causal connections among the six regions should not exist if there is no anatomical connection. To obtain the final estimated graph, note that in the computing MPC, we have solved a linear sum assignment problem, which matches the semantic information of the estimated variables with the one of the true underlying latent variables. After matching all semantic information, we can use a threshold (0.1) to removed a edge separately across $\mathbf{u}$, and once an edge appears in more than $70\%$ of all $\mathbf{u}$, we keep this edge in the final graph. Figure 19 (c) shows an example of the estimated causal skeleton by the proposed SuaVE. Averaging over 5 different random seeds for the used nonlinear mapping from latent space to observed space, structural Hamming distance (SHD) obtained by the proposed SuaVE is 5.0 $\pm$ 0.28. In contrast, iVAE, $\beta$-VAE and VAE assume latent variables to be independent, and thus can not obtain the causal skeleton. We further analyse the result of $\beta$-VAE with $\beta=4$ by using the PC algorithm (Spirtes et al., 2001) to discovery skeleton on the recovered latent variables, and obtain SHD = 5.8 $\pm$ 0.91. This means that: 1) $\beta$-VAE with $\beta=4$ does not ensure the strong independence among the recover latent variables as it is expected, 2) the proposed SuaVE is an effective one-step method to discovery the skeleton. The unsupervised version of CausalVAE is not identifiable, and can also not ensure the relations among the recover variables to be causal skeleton in principle. In experiments we found that CausalVAE always obtains fully-connected graphs for the 5 different random seeds, for which SHD is 9.0 $\pm$ 0.
Figure 18: Performances of $\beta$-VAE with different $\beta$ in recovering fMRI data.
We further investigate the performance of $\beta$-VAE across different values of $\beta$. Figure 18 illustrates the performance trend as $\beta$ varies. $\beta$-VAE suffers from two critical limitations: In practical applications, determining an appropriate value for $\beta$ is challenging due to the inherent uncertainty in the underlying latent variables. This lack of guidance can lead to suboptimal model performance if $\beta$ is not carefully tuned. For $\beta\geq 1$, as the value of $\beta$ increases, there is an inevitable trade-off between disentanglement and reconstruction quality. Specifically, improving disentanglement typically comes at the cost of degraded reconstruction, as compared to a standard VAE. This trade-off is particularly problematic in scenarios where preserving reconstruction quality is crucial, making it difficult to balance these competing objectives in real-world tasks. These challenges highlight the practical difficulties in applying $\beta$-VAE effectively, especially when considering applications where the precise nature of the latent structure is unknown and both disentanglement and reconstruction are essential.
Figure 19: Performance on fMRI Data. (a) the performance of the proposed SuaVE in comparison to iVAE, $\beta$-VAE, VAE and CausalVAE in recovering the latent six signals. (b) The skeleton of the anatomical connections given in Bird and Burgess (2008). (c) The recovered skeleton by the proposed SuaVE, where the dashed lines indicate errors.
Figure 20: Recovered graphs on fMRI Data. (a) the anatomical connections given in Bird and Burgess (2008). (b) The recovered causal graph by the proposed SuaVE. (c) The recovered causal graph reported in (Ghassami et al., 2018). Note that the setting in (b) treats fMRI data as latent variables, while the setting in (c) treats fMRI data as observed variables.
Figure 20 (a) show the edges in the anatomical ground truth reported in (Ghassami et al., 2018). Figure 20 (b) show the latent recovered edges by the proposed SuaVE, in the setting where we treat fMRI data as latent causal variables. For comparison, Figure 20 (c) shows the recovered causal graph reported in (Ghassami et al., 2018), in the setting where fMRI data are treated as observed variables. That is, the setting in (b) is more challenging than the setting in (c).
8 Conclusion
Identifying latent causal representations is known to be generally impossible without certain assumptions. Motivated by recent progress in the identifiability result of nonlinear ICA, we analyse three intrinsic indeterminacies in latent space, e.g., transitivity, permutation indeterminacy and scaling indeterminacy, which provide deep insights for identifying latent causal representations. To address the unidentifiable issue due to transitivity, we explore a new line of research in the identifiability of latent causal representations, that of allowing the causal influences among the latent causal variables to change. We show that latent causal representations, modeled by weight-variant linear Gaussian models, can be identified up to permutation and scaling, with mild assumptions. We then propose a novel method to consistently learn latent causal representations. Experiments on synthetic and fMRI data demonstrate the identifiability and consistency results and the efficacy of the proposed method.
Acknowledgments and Disclosure of Funding
Appendix A
A.1 The proof of Theorem 1
The proof of Theorem 1 is done in three steps. Step I is to show the identifiability result of nonlinear ICA holds in our setting where there is an additional mapping from $\mathbf{n}$ to $\mathbf{z}$, and the mapping changes across $\mathbf{u}$. Step II consists of building a linear transformation between the true latent variables $\mathbf{z}$ and the estimated ones $\mathbf{\hat{z}}$. Step III shows that the linear transformation in Step II can be reduced to permutation transformation by using the result of Step I and the assumption (v).
For convenience, let us discuss the property of the mapping from $\mathbf{n}$ to $\mathbf{z}$ first, which plays a key role in the proof. Let function $\mathbf{g}_{\mathbf{u}}$ denote the mapping from $\mathbf{n}$ to $\mathbf{z}$. That is:
$\mathbf{z}=\mathbf{g}_{\mathbf{u}}(\mathbf{n})=(\mathbf{I}-\boldsymbol{\lambda}^{T}(\mathbf{u}))^{-1}\mathbf{n},$ (14)
where $\mathbf{I}$ denotes the identity matrix. Due to DAG constraints in a causal system and the assumption of linear models on $\mathbf{z}$, we have 1) the determinant of the Jacobian matrix is equal to 1, e.g., $|\det{\mathbf{J}}_{\mathbf{g_{u}}^{-1}}(\mathbf{z})|=|\det{\mathbf{J}}_{\mathbf{g_{u}}}(\mathbf{u})|=1$, 2) the mapping $\mathbf{g}_{\mathbf{n}}$ is invertible.
Step I: Suppose we have two sets of parameters $\theta=({\mathbf{f},\mathbf{T}_{\mathbf{n}},\boldsymbol{\lambda},\boldsymbol{\beta}},\boldsymbol{\varepsilon})$ and $\hat{\theta}=({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{n}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})$ such that $p_{({\mathbf{f},\mathbf{T}_{\mathbf{n}},\boldsymbol{\lambda},\boldsymbol{\beta}},\boldsymbol{\varepsilon})}(\mathbf{x}|\mathbf{u})=p_{({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{n}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})}(\mathbf{x}|\mathbf{u})$ for all pairs $(\mathbf{x},\mathbf{u})$. Due to the assumptions (i) and (ii), by expanding these expressions via the change of variables formula and taking the logarithm we find:
$\log|{\mathbf{J}_{{\mathbf{f}}^{-1}}(\mathbf{x})}|+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})+$ $\log|{\mathbf{J}_{\mathbf{g}^{-1}_{\mathbf{u}}}(\mathbf{z})}|+\log p_{\mathbf{(\mathbf{T}_{\mathbf{n}},\boldsymbol{\beta})}}({\mathbf{n}|\mathbf{u}})$
$=\log{|\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\log{|\mathbf{J}_{(\mathbf{\hat{g}}_{\mathbf{u}})^{-1}}(\mathbf{\hat{z}})|}+\log p_{\mathbf{(\mathbf{\hat{T}}_{\mathbf{n}},\boldsymbol{\hat{\beta}})}}({\mathbf{\hat{n}}|\mathbf{u}}),$ (15)
Using the exponential family Eq. (4) to replace $p_{\mathbf{(\mathbf{T}_{\mathbf{n}},\boldsymbol{\beta})}}({\mathbf{n}|\mathbf{u}})$, we have:
$\log|{\mathbf{J}_{{\mathbf{f}}^{-1}}(\mathbf{x})}|+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})+\log|{\mathbf{J}_{\mathbf{g}^{-1}_{\mathbf{u}}}(\mathbf{z})}|+\mathbf{T}^{T}_{\mathbf{n}}\big{(}\mathbf{n}\big{)}\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u})-\log{Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u})}=$ (16)
$\log{|\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\log{|\mathbf{J}_{(\mathbf{\hat{g}}_{\mathbf{u}})^{-1}}(\mathbf{\hat{z}})|}+\mathbf{\hat{T}}_{\mathbf{n}}^{T}\big{(}\mathbf{\hat{n}}\big{)}\boldsymbol{\hat{\eta}}_{\mathbf{n}}(\mathbf{u})-\log{\hat{Z}_{\mathbf{n}}(\boldsymbol{\hat{\beta}},\mathbf{u})},$ (17)
then by using
$|\det{\mathbf{J}}_{\mathbf{g_{u}}^{-1}}(\mathbf{z})|=1,|\det{\mathbf{J}}_{\mathbf{\hat{g}_{u}}^{-1}}(\mathbf{z})|=1.$ (18)
here the latter holds since $\hat{g}_{u}$ must be the same function class as $g_{u}$, Eqs. 16 and 17 can be rewriteed as:
$\log{|\mathbf{J}_{\mathbf{f}^{-1}}(\mathbf{x})|}+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})+\mathbf{T}^{T}_{\mathbf{n}}\big{(}\mathbf{n}\big{)}\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u})-\log{Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u})}=$ (19)
$\log{|\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\mathbf{\hat{T}}_{\mathbf{n}}^{T}\big{(}\mathbf{\hat{n}}\big{)}\boldsymbol{\hat{\eta}}_{\mathbf{n}}(\mathbf{u})-\log{\hat{Z}_{\mathbf{n}}(\boldsymbol{\hat{\beta}},\mathbf{u})}.$ (20)
The following proof is similar to the proof of Sorrenson et al. (2020). For completeness, we present a slightly simplified proof. With different points in assumption (iii), we can subtract this expression with $\mathbf{u}_{\mathbf{n},0}$ by the expression with $l_{\mathbf{n}}$, $l_{\mathbf{n}}=1,...,2n$. Since the Jacobian and noise terms do not depend on $\mathbf{u}$, the Jacobian and noise terms will be canceled out:
$\log\frac{Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u}_{\mathbf{n},0})}{Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u}_{\mathbf{n},l_{\mathbf{n}}})}+\mathbf{T}^{T}_{\mathbf{n}}\big{(}\mathbf{n}\big{)}\big{(}\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},l_{\mathbf{n}}})-\boldsymbol{\eta}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},0})\big{)}=$ (21)
$\log\frac{\hat{Z}_{\mathbf{n}}(\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{n},0})}{\hat{Z}_{\mathbf{n}}(\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{n},l_{\mathbf{n}}})}+\mathbf{\hat{T}}_{\mathbf{n}}^{T}\big{(}\mathbf{\hat{n}}\big{)}\big{(}\boldsymbol{\hat{\eta}}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},l_{\mathbf{n}}})-\boldsymbol{\hat{\eta}}_{\mathbf{n}}(\mathbf{u}_{\mathbf{n},0})\big{)}.$ (22)
Combining the $2n$ expressions into a single matrix equation we can write it in terms of $\mathbf{L}_{\mathbf{n}}$ from assumption (iii):
$\mathbf{L}^{T}_{\mathbf{n}}\mathbf{T_{\mathbf{n}}\big{(}\mathbf{n}\big{)}}=\mathbf{\hat{L}}^{T}_{\mathbf{n}}\mathbf{\hat{T}_{\mathbf{n}}\big{(}\mathbf{\hat{n}}\big{)}}+\mathbf{b}_{\mathbf{n}},$ (23)
where $\mathbf{L}_{\mathbf{n}}$ is defined as the assumption (iii), $\mathbf{b}_{\mathbf{n},l_{\mathbf{n}}}=\log\frac{\hat{Z}_{\mathbf{n}}(\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{n},0})Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u}_{\mathbf{n},l_{\mathbf{n}}})}{\hat{Z}_{\mathbf{n}}(\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{n},l_{\mathbf{n}}})Z_{\mathbf{n}}(\boldsymbol{\beta},\mathbf{u}_{\mathbf{n},0})}$. Since $\mathbf{L}^{T}_{\mathbf{n}}$ is invertible by the assumption (iii), we can multiply this expression by its inverse:
$\mathbf{T_{\mathbf{n}}\big{(}\mathbf{n}\big{)}}=\mathbf{A}_{\mathbf{n}}\mathbf{\hat{T}_{\mathbf{n}}\big{(}\mathbf{\hat{n}}\big{)}}+\mathbf{c}_{\mathbf{n}},$ (24)
According to lemma 3 in (Khemakhem et al., 2020) and Proof for Theorem 1 IN (Sorrenson et al., 2020) we can show that $\mathbf{A}_{\mathbf{n}}$ has full rank. Since we assume the noise to be Gaussian, Eq. (24) can be re-expressed as:
$\left(\begin{array}[]{c}\mathbf{n}\\ \mathbf{n}^{2}\\ \end{array}\right)=\mathbf{A}_{\mathbf{n}}\left(\begin{array}[]{c}\mathbf{\hat{n}}\\ \mathbf{\hat{n}}^{2}\\ \end{array}\right)+\mathbf{c}_{\mathbf{n}},$ (25)
Then, by the contradiction between 1) for every $n_{i}$, there exists a polynomial with a degree at most 2, and 2) for every $n^{2}_{i}$, there also exists a polynomial with a degree at most 2 (Sorrenson et al., 2020), we can obtain:
$\mathbf{n}=\mathbf{P}\mathbf{\hat{n}}+\mathbf{c}_{\mathbf{n}},$ (26)
where $\mathbf{P}$ denote the permutation matrix with scaling. More details can be found in (Sorrenson et al., 2020).
Step II: Again, suppose we have two sets of parameters $\theta=({\mathbf{f},\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta}})$ and $\hat{\theta}=({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{z}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})$ such that $p_{({\mathbf{f},\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta}})}(\mathbf{x}|\mathbf{u})=p_{({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{z}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})}(\mathbf{x}|\mathbf{u})$ for all pairs $(\mathbf{x},\mathbf{u})$. Due to the assumptions (i) and (ii), by expanding these expressions via the change of variables formula and taking the logarithm we have:
$\log{|\det\mathbf{J}_{\mathbf{f}^{-1}}(\mathbf{x})|}+\log p_{\mathbf{(\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta})}}({\mathbf{z}|\mathbf{u}})+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})=\log{|\det\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\log p_{\mathbf{(\mathbf{\hat{T}}_{\mathbf{z}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}})}}({\mathbf{\hat{z}}|\mathbf{u}}),$ (27)
Using the exponential family Eq. (7) to replace $p_{\mathbf{(\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta})}}({\mathbf{z}|\mathbf{u}})$:
$\log{|\det\mathbf{J}_{\mathbf{f}^{-1}}(\mathbf{x})|}+\mathbf{T}^{T}_{\mathbf{z}}(\mathbf{z})\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u})-\log{Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u})}+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})=$ (28)
$\log{|\det\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\mathbf{\hat{T}}^{T}_{\mathbf{z}}(\mathbf{\hat{z}})\boldsymbol{\hat{\eta}}_{\mathbf{z}}(\mathbf{u})-\log{\hat{Z}_{\mathbf{z}}(\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}},\mathbf{u})}.$ (29)
Let $\mathbf{u}_{\mathbf{z},0},\mathbf{u}_{\mathbf{z},1},...,\mathbf{u}_{\mathbf{z},k}$ be the points provided by the assumption (iv). We define $\bar{\boldsymbol{\eta}}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})={\boldsymbol{\eta}}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})-{\boldsymbol{\eta}_{\mathbf{z}}}(\mathbf{u}_{\mathbf{z},0})$. We plug each of those $\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}},l_{\mathbf{z}}=1,...,k$ in Eq. (29) to obtain $k+1$ equations. We subtract the first equation for $\mathbf{u}_{\mathbf{z},0}$ from the remaining $k$ equations. For example, for $\mathbf{u}_{\mathbf{z},0}$ and $\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}}$ we have two equations:
$\log{|\det\mathbf{J}_{\mathbf{f}^{-1}}(\mathbf{x})|}+\mathbf{T}^{T}_{\mathbf{z}}(\mathbf{z})\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},0})-\log{Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u}_{\mathbf{z},0})}+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})=$ (30a)
$\log{|\det\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\mathbf{\hat{T}}^{T}_{\mathbf{z}}(\mathbf{\hat{z}})\boldsymbol{\hat{\eta}}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},0})-\log{\hat{Z}_{\mathbf{z}}(\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{z},0})},$
$\log{|\det\mathbf{J}_{\mathbf{f}^{-1}}(\mathbf{x})|}+\mathbf{T}^{T}_{\mathbf{z}}(\mathbf{z})\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})-\log{Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})}+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})=$ (30b)
$\log{|\det\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\mathbf{\hat{T}}^{T}_{\mathbf{z}}(\mathbf{\hat{z}})\boldsymbol{\hat{\eta}}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})-\log{\hat{Z}_{\mathbf{z}}(\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})},$
Using Eq. (30b) subtracts Eq. (30a), canceling the terms that do not include $\mathbf{u}$, we have:
$\log\frac{Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u}_{\mathbf{z},0})}{Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})}+\mathbf{T}_{\mathbf{z}}{(\mathbf{z})}\big{(}\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},l_{\mathbf{n}}})-\boldsymbol{\eta}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},0})\big{)}$ (31)
$\qquad\qquad\qquad=\log\frac{\hat{Z}_{\mathbf{z}}(\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{z},0})}{\hat{Z}_{\mathbf{z}}(\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})}+\mathbf{\hat{T}}_{\mathbf{z}}{(\mathbf{\hat{z}})}\big{(}\boldsymbol{\hat{\eta}}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},l_{\mathbf{n}}})-\boldsymbol{\hat{\eta}}_{\mathbf{z}}(\mathbf{u}_{\mathbf{z},0})\big{)}.$
Let $\mathbf{L}_{\mathbf{z}}$ be the matrix defined in Eq. (9) in assumption (iv), and $\mathbf{\hat{L}}_{\mathbf{\hat{z}}}$ similarly ($\mathbf{\hat{L}}_{\mathbf{\hat{z}}}$ is not necessarily invertible). Define $b_{\mathbf{z},l_{\mathbf{z}}}={\log{\frac{\hat{Z}_{\mathbf{z}}(\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{z},0})Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})}{Z_{\mathbf{z}}(\boldsymbol{\lambda},\boldsymbol{\beta},\mathbf{u}_{\mathbf{z},0})\hat{Z}_{\mathbf{z}}(\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}},\mathbf{u}_{\mathbf{z},l_{\mathbf{z}}})}}}$. Expressing Eq. (31) for all points in matrix form, we get:
$\mathbf{L}_{\mathbf{z}}^{T}\mathbf{T_{\mathbf{z}}(\mathbf{z})}=\mathbf{\hat{L}}_{\mathbf{z}}^{T}\mathbf{\hat{T}_{\mathbf{z}}(\mathbf{\hat{z}})}+\mathbf{b}_{\mathbf{z}}.$ (32)
We multiply both sides of Eq. (32) by the inverse of $\mathbf{L}^{T}_{\mathbf{z}}$ (by assumption (iv)) to find:
$\mathbf{T_{\mathbf{z}}(\mathbf{z})}=\mathbf{A}_{\mathbf{z}}\mathbf{\hat{T}_{\mathbf{z}}(\mathbf{\hat{z}})}+\mathbf{c}_{\mathbf{z}},$ (33)
where $\mathbf{A}_{\mathbf{z}}=(\mathbf{L}_{\mathbf{z}}^{T})^{-1}\mathbf{\hat{L}}_{\mathbf{\hat{z}}}^{T}$ and $\mathbf{c}_{\mathbf{z}}=(\mathbf{L}_{\mathbf{z}}^{T})^{-1}\mathbf{b}_{\mathbf{z}}$. As mentioned in Eq. (7), the sufficient statistic $\mathbf{T}_{\mathbf{z}}(\mathbf{z})=[\mathbf{z};\operatorname{vec}(\mathbf{z}\mathbf{z}^{T})]$. In this case, the relationship Eq. (33) becomes:
$\left(\begin{array}[]{c}\mathbf{z}\\ \mathbf{z}^{2}\\ \mathbf{z}_{i\neq j}\\ \end{array}\right)=\mathbf{A}_{\mathbf{z}}\left(\begin{array}[]{c}\mathbf{\hat{z}}\\ \mathbf{\hat{z}}^{2}\\ \mathbf{\hat{z}}_{i\neq j}\\ \end{array}\right)+\mathbf{c}_{\mathbf{z}},$ (34)
where $\mathbf{z}$ denotes $[z_{1},...,z_{i}]$, $\mathbf{z}^{2}$ denotes $[z^{2}_{1},...,z^{2}_{i}]$, $\mathbf{z}_{i\neq j}$ denotes the vector whose elements are ${z}_{i}{z}_{j}$ for all $i\neq j$, $\mathbf{A}_{\mathbf{z}}$ in block matrix can be rewritten as:
$\mathbf{A}_{\mathbf{z}}=\left(\begin{array}[]{ccc}\mathbf{A}^{(1)}&\mathbf{A}^{(2)}&\mathbf{A}^{(3)}\\ \mathbf{A}^{(4)}&\mathbf{A}^{(5)}&\mathbf{A}^{(6)}\\ \mathbf{A}^{(7)}&\mathbf{A}^{(8)}&\mathbf{A}^{(9)}\\ \end{array}\right)$ (35)
and $\mathbf{c}_{\mathbf{z}}$ as:
$\mathbf{c}_{\mathbf{z}}=\left(\begin{array}[]{c}\mathbf{c}^{(1)}\\ \mathbf{c}^{(2)}\\ \mathbf{c}^{(3)}\\ \end{array}\right).$ (36)
Then, we have:
$\mathbf{z}=\mathbf{A}^{(1)}\mathbf{\hat{z}}+\mathbf{A}^{(2)}\mathbf{\hat{z}}^{2}+\mathbf{A}^{(3)}\mathbf{\hat{z}}_{i\neq j}+\mathbf{c}^{(1)},$ (37)
$\mathbf{z}^{2}=\mathbf{A}^{(4)}\mathbf{\hat{z}}+\mathbf{A}^{(5)}\mathbf{\hat{z}}^{2}+\mathbf{A}^{(6)}\mathbf{\hat{z}}_{i\neq j}+\mathbf{c}^{(2)}.$ (38)
So we can write for each ${z}_{i}$:
${z_{i}}=\sum_{j}({A}_{i,j}^{(1)}{\hat{z}}_{j})+\sum_{j}({A}_{i,j}^{(2)}{\hat{z}^{2}}_{j})+({\mathbf{A}}_{i,:}^{(3)}\hat{\mathbf{z}}_{i\neq j})+{c}_{i}^{(1)},$ (39)
$z_{i}^{2}=\sum_{j}({A}_{i,j}^{(4)}{\hat{z}}_{j})+\sum_{j}({A}_{i,j}^{(5)}{\hat{z}^{2}}_{j})+({\mathbf{A}}_{i,:}^{(6)}\hat{\mathbf{z}}_{i\neq j})+{c}_{i}^{(2)}.$ (40)
Squaring Eq. (39), we have:
${z}_{i}^{2}=$ $\underbrace{(\sum_{j}({A}_{i,j}^{(2)}{\hat{z}^{2}}_{j}))^{2}}_{\mathbf{(a)}}+\underbrace{(\sum_{j}({A}_{i,j}^{(1)}{\hat{z}}_{j}))^{2}}_{\text{(b)}}+\underbrace{({\mathbf{A}}_{i,:}^{(3)}\hat{\mathbf{z}}_{i\neq j})^{2}}_{\text{(c)}}+({c}_{i}^{(1)})^{2}+.....$ (41)
It is notable that Eq. (40) and Eq. (41) are derived from Eq. (33) which holds for arbitrary $\mathbf{x}$, then by the assumption (ii) in Theorem 1 that $\mathbf{f}$ is a bijective mapping from $\mathbf{z}$ to $\mathbf{x}$, and thus Eq. (40) and Eq. (41) must holds for $\mathbf{z}$ everywhere. Then by the fact that the right sides of the Eq. (40) and Eq. (41) are both polynomials with finite degree, we have each coefficients of the two polynomials must be equal. In more detail, for the term (a) in Eq. (41):
${(\sum_{j}({A}_{i,j}^{(2)}{\hat{z}^{2}}_{j}))^{2}}=\sum_{j}({A}_{i,j}^{(2)})^{2}{\hat{z}^{4}}_{j}+\sum_{j\neq j^{\prime}}(2{A}_{i,j}^{(2)}{A}_{i,j^{\prime}}^{(2)}{\hat{z}^{2}}_{j}{\hat{z}^{2}}_{j^{\prime}}).$ (42)
Compared with Eq. (40), since there is no term ${\hat{z}^{4}}_{j}$ in Eq. (40), we must have that:
$\mathbf{A}^{(2)}=0.$ (43)
For the term (c) in Eq. (41):
${({\mathbf{A}}_{i,:}^{(3)}\hat{\mathbf{z}}_{i\neq j})^{2}}={\sum_{j^{{}^{\prime}},i\neq j}({A}_{i,j^{\prime}}^{(3)})^{2}(\hat{z}_{i}\hat{z}_{j})^{2}}+...$ (44)
Compared with Eq. (40), since there is no term $(\hat{z}_{i}\hat{z}_{j})^{2}$ in Eq. (40), we must have that:
$\mathbf{A}^{(3)}=0$ (45)
As a result, Eq. (37) becomes:
$\mathbf{z}=\mathbf{A}^{(1)}\mathbf{\hat{z}}+\mathbf{c}^{(1)}.$ (46)
The above equation indicates the latent causal variables can be recovered up to linear transformation.
Step III We then show that the linear transformation matrix $\mathbf{A}^{(1)}$ in Eq. (46) must be a permutation matrix. As we mentioned in step I, for the Gaussian noise variables $\mathbf{n}$, we have:
$\mathbf{n}=\mathbf{P}\mathbf{\hat{n}}+\mathbf{c}_{n},$ (47)
where $\mathbf{P}$ denote the permutation matrix with scaling. With Eq. (47) and Eq. (14), the Eq. (46) can be rewritten as follows:
$\mathbf{B}(\mathbf{P}\mathbf{\hat{n}}+\mathbf{c}_{n})=\mathbf{A}^{(1)}(\mathbf{\hat{B}}\mathbf{\hat{n}})+\mathbf{c}^{(1)}.$ (48)
$\Rightarrow$ $(\mathbf{B}\mathbf{P}-\mathbf{A}^{(1)}\mathbf{\hat{B}})\mathbf{\hat{n}}=\mathbf{c}^{(1)}-\mathbf{B}\mathbf{c}_{n},$ (49)
where $\mathbf{B}=(\mathbf{I}-\boldsymbol{\lambda}^{T}(\mathbf{u}))^{-1}$ and $\mathbf{\hat{B}}=(\mathbf{I}-\boldsymbol{\hat{\lambda}}^{T}(\mathbf{u}))^{-1}$. By differentiating Eq. (49) with respect to $\mathbf{\hat{n}}$, we have:
$(\mathbf{B}\mathbf{P}-\mathbf{A}^{(1)}\mathbf{\hat{B}})\mathbf{I}=\mathbf{0},$ (50)
where $\mathbf{I}$ denote the identity matrix, which implies that:
$\mathbf{B}\mathbf{P}=\mathbf{A}^{(1)}\mathbf{\hat{B}},$ (51)
that is,
$(\mathbf{I}-\boldsymbol{\lambda}^{T}(\mathbf{u}))^{-1}\mathbf{P}=\mathbf{A}(\mathbf{I}-\boldsymbol{\hat{\lambda}}^{T}(\mathbf{u}))^{-1},$ (52)
for convenience, we use $\mathbf{A}$ to replace $\mathbf{A}^{(1)}$ here and in the following. By inverting both sides of the equation, we have
$\mathbf{P}^{-1}(\mathbf{I}-\boldsymbol{\lambda}^{T}(\mathbf{u}))=(\mathbf{I}-\boldsymbol{\hat{\lambda}}^{T}(\mathbf{u}))\mathbf{A}^{-1}.$ (53)
Without loss of generality, suppose that $\boldsymbol{\lambda}^{T}$ and $\boldsymbol{\hat{\lambda}}^{T}$ denote lower triangular matrices, respectively.
Elements above the diagonal
We can see from Eq. (53) that since the left is a lower triangular matrix, and $(\mathbf{I}-\boldsymbol{\hat{\lambda}}^{T}(\mathbf{u}))$ in the right is also a lower triangular matrix, $\mathbf{A}^{-1}$ must be a lower triangular matrix. That is, elements above the diagonal in matrix $\mathbf{A}^{-1}$ are equal to 0.
The diagonal elements
Assume $\mathbf{P}^{-1}$ to be diagonal with elements $s_{1,1},s_{2,2},s_{3,3},...$. By expanding the both sides of Eq. (53), we have:
$\left(\begin{array}[]{cccc}s_{1,1}&0&0&...\\ -s_{2,2}\lambda_{1,2}(\mathbf{u})&s_{2,2}&0&...\\ -s_{3,3}\lambda_{1,3}(\mathbf{u})&-s_{3,3}\lambda_{2,3}(\mathbf{u})&s_{3,3}&...\\ .&.&.&...\\ \end{array}\right)$ (54)
$=\left(\begin{array}[]{cccc}A_{1,1}&0&0&...\\ A_{2,1}-A_{1,1}\hat{\lambda}_{1,2}(\mathbf{u})&A_{2,2}&0&...\\ A_{3,1}-\sum_{i=1}^{2}A_{i,1}\hat{\lambda}_{i,3}(\mathbf{u})&A_{3,2}-A_{2,2}\hat{\lambda}_{2,3}(\mathbf{u})&A_{3,3}&...\\ .&.&.&...\\ \end{array}\right).$
Comparing the left and the right of Eq. (54), we can see that the diagonal elements of $\mathbf{A}^{-1}$ must be $s_{1,1},s_{2,2},s_{3,3},...$.
Elements below the diagonal
By comparison between the left and right of (54), for each element ${A}_{i,j}$ where $i>j$ in the right of Eq. (54), we have: ${A}_{i,j}-\sum_{i^{\prime}<i}{\mathbf{A}_{i^{\prime},j}{\hat{\lambda}_{i^{\prime},j}}(\mathbf{u})}=s_{j,j}\lambda_{j,i}(\mathbf{u})$. Then by assumption (v), we have ${A}_{i,j}=0$.
As a result, the matrix $\mathbf{A}^{-1}$ is the same as the matrix $\mathbf{P}^{-1}$. That is,
$\mathbf{z}=\mathbf{P}\mathbf{\hat{z}}+\mathbf{c}_{\mathbf{z}}.$ (55)
A.2 Understanding Assumptions in Theorem 4.1
Gaussian assumption on the latent noise variables
Our first assumption is implicitly enforcing Gaussian distribution on the latent noise variables. Note that the assumptions of nonlinear ICA (Khemakhem et al., 2020) on the noise could be broad exponential family distribution, e.g., Gaussian distributions, Laplace, Gamma distribution, and so on. This work considers Gaussian distribution, mainly because it can be straightforwardly implemented for the re-parameterization trick in VAE. 2) it is convenient for simplifying the proof process. It is worthwhile and promising to extend the Gaussian assumption to exponential family for future work.
Linear model assumption for the latent causal variables
As the first work to discuss the relation between the change of causal influences and identifiability of latent causal model, this work simply considers linear models for the latent causal variables, because it can be directly parameterized, which helps us to analyze the challenge of identifiability and how to handle it. In addition, together with Gaussian noise, we arrive at the same probabilistic model as linear Gaussian Bayesian networks, which is a well-studied model as mentioned in Eq. (6). We expect that the proposed weights-variant linear models can motivate more general functional classes in future work, e.g., nonlinear additive noise models.
Changes of weights
We allow changes of weight (e.g., causal influences) among latent causal variables across $\mathbf{u}$ to handle the transitivity problem. We argue that this is not a necessary condition for identifiability (Existing methods, including sparse graph structure and temporal information, could be regarded as feasible ways to handle the transitivity.), but it is a sufficient condition to handle the transitivity, and provide a new research line for causal representation learning. In fact, exploring the change of causal influences is not new in observed space (Ghassami et al., 2018; Huang et al., 2020). In addition, identifying latent causal representation by randomly chosen unknown hard intervention Brehmer et al. (2022) can also be regarded as a special change of causal influences. For this viewpoint, we argue that exploring the change of causal influences for identifying latent causal representation is a promising way and not too restricted in reality.
We emphasize that we require the change of weights among latent causal variables across $\mathbf{u}$. But this does not imply that the weights always change across $\mathbf{u}$. That is, some weights are allowed to keep unchanged for some points $\mathbf{u}$, as long as these weights change for other points $\mathbf{u}^{\prime}$. See experimental results in section 7.1.1 for further discussion.
A.3 The Proof of Corollary 2
Theorem 1 has shown that the latent causal variables $\mathbf{z}$ can be identified up to trivial permutation and linear scaling. Hence, the identifiability of the causal structure in latent space can be reduced to the identifiability of the causal structure in observed space. Moreover, we need to show that the linear scaling does not affect theoretical identifiability of the causal structure.
For the identifiability of the causal structure in observed space from heterogeneous data, fortunately, we can leverage the results from Huang et al. (2020). Corollary 4.2 relies on the Markov condition and faithfulness assumption and the assumption that the latent change factor (i.e., causal strength in the linear case) can be represented as a function of the domain index $\mathbf{u}$. Hence, it relies on the same assumptions as that in Huang et al. (2020).
It has been shown in Huang et al. (2020) that if the joint distribution over $\mathbf{z}$ and $\mathbf{u}$ ($\mathbf{z}$ and $\mathbf{u}$ are observed variables here) are Markov and faithful to the augmented graph, then the causal structure over $\mathbf{z}\cup\mathbf{u}$ can be identified up to the Markov equivalence class, by making use of the conditional independence relationships.
Next, we show that the Markov equivalence class over $\mathbf{z}$ is also identifiable. Denote by $M$ the Markov equivalence class over $\mathbf{z}\cup\mathbf{u}$, and by $M_{z}$ the Markov equivalence class over $\mathbf{z}$. Then after removing variable $\mathbf{u}$ in $M$ and its edges, the resulting graph (denoted by $M’_{z}$) is the same as $M_{z}$. This is because of the following reasons. First, it is obvious that $M’_{z}$ and $M_{z}$ have the same skeleton. Second, in this paper, $\mathbf{u}$ has an edge over every $z_{i}$ when considering $\mathbf{n}$ as latent noise variables, because all causal strength and noise distributions change with $\mathbf{u}$. Hence, there is no v-structure over $\mathbf{u}$, $z_{i}$, and $z_{j}$, so it is not possible to have more oriented edges in $M’_{z}$. Therefore, $M’_{z}$ and $M_{z}$ have the same skeleton and the same directions.
Moreover, the conditional independence relationships will not be affected by the linear scaling of the variables, so the conditional independence relationships still hold in the identified latent variables. Furthermore, for linear-Gaussian models, independence equivalence is the same as distributional equivalence. This is because of the following reasons. First, since we are concerned with linear Gaussian models over the latent variables, the identifiability up to equivalence class also holds for score-based methods that use the likelihood of data as objective functions. This is because independence equivalence, i.e., two DAGs have the same conditional independence relations, is the same as distributional equivalence for linear-Gaussian models, i.e., two Bayesian networks corresponding to the two DAGs can define the same probability distribution. That is to say, score-based methods also find the structure based on independence relations implicitly. Because the scaling does not change the independence relations, it will also not affect the identifiability of the graph structure.
Therefore, the causal structure among latent variables $\mathbf{z}$ can be identified up to the Markov equivalence class.
A.4 The Proof of Corollary 3
To prove this corollary, let us introduce ICM first.
Definition 5
(Independent Causal Mechanisms). In a causally sufficient system, the causal modules, as well as their included parameters, change independently across domains.
Here we follow Ghassami et al. (2018) to define ICM principle. Another slightly different definition appears in Peters et al. (2017); Schölkopf et al. (2021). Both two definitions imply the same principle that the causal modules change independently across $\mathbf{u}$. According to ICM principle, in the causal direction, the causal modules, as well as their included parameters, change independently across $\mathbf{u}$. However, such independence generally does not hold in the anti-causal direction (Ghassami et al., 2018; Huang et al., 2020). For example, consider $z_{1}$ and $z_{2}$ in Figure 1. According to model definition in Eq. (1)- Eq. (3), for causal direction $z_{1}\rightarrow z_{2}$, we have:
$n_{i}:\sim{\cal N}({\beta_{i,1}(\mathbf{u})},{\beta_{i,2}(\mathbf{u})}),\qquad z_{1}:=n_{1},\qquad z_{2}:=\lambda_{1,2}({\mathbf{u}})z_{1}+n_{2}.$ (56)
For the reverse direction, we have
$\begin{split}&z_{1}=\lambda^{\prime}_{1}(\mathbf{u})z_{2}+n_{1}^{\prime},\qquad\qquad\qquad\qquad\ z_{2}=n_{2}^{\prime}\\ &\lambda^{\prime}_{1}(\mathbf{u})=\frac{1}{\lambda_{1,2}(\mathbf{u})},\quad n_{1}^{\prime}=-\frac{1}{\lambda_{1,2}(\mathbf{u})}n_{2}\qquad n_{2}^{\prime}=\lambda_{1,2}({\mathbf{u}})z_{1}+n_{2}\end{split}$ (57)
In this case, the module $p(z_{2})$ is dependent of the module $p(z_{1}|z_{2})$, e.g., both modules depend on the same parameter $\lambda_{1,2}({\mathbf{u}})$, which violates ICM principle. Clearly, the scaling indeterminacy for the recovered latent variables does not affect such dependence. As a result, regardless of the scaling, with the help of ICM, we can fully identify the acyclic causal structure among latent variables $\mathbf{z}$.
A.5 The proof of Theorem 4
Throughout the proof process in Section A.1, Eq. (54) holds without the need for assumption (v). Note that this also implies that $\mathbf{A}^{-1}$ is a lower triangular matrix. Then consider the following two cases.
  • For the case where $z_{i}$ is a root node or all weights on all paths from parent nodes of $z_{i}$ to $z_{i}$ meet assumption (v), by using assumption (v) and by comparison both sides of Eq. (54), we have: for all $j<i$, ${A}_{{i,j}}=0$, which implies that we can obtain that $A^{-1}_{i,i}z_{i}=\hat{z}_{i}+c^{\prime}_{i}$.
  • If there exists an unchanged weight on all paths from parent nodes of $z_{i}$ to $z_{i}$ across $\mathbf{u}$, we demonstrate that it is always possible to construct an alternative solution, which is different from the true $\mathbf{z}$, but capable of generating the same observations $\mathbf{x}$. Suppose that for $z_{i}$, there is an unchanged weight $\lambda_{j,i}$ across $\mathbf{u}$, related to the parent node $z_{j}$. Then, we can always construct new latent variables $\mathbf{z^{\prime}}$ as: for all $k\neq i$, ${z^{\prime}}_{k}=z_{k}$, and $z^{\prime}_{i}=z_{i}-\lambda_{j,i}z_{j}$. Given this, we can construct a matrix $\mathbf{M}$, so that $\mathbf{M}\mathbf{z^{\prime}}=\mathbf{z}$. It is clear the matrix $\mathbf{M}$ is invertible. In addition, all the weights of the matrix $\mathbf{M}$ are constant and thus do not depend on $\mathbf{u}$. As a result, we can construct a mapping from $\mathbf{z}^{\prime}$ to $\mathbf{x}$, e.g., $\mathbf{f}(\mathbf{M}\mathbf{z}^{\prime})$, which is invertible and do not depend on $\mathbf{u}$, and can create the same data $\mathbf{x}$ generated by $\mathbf{f}(\mathbf{z})$. Therefore, the alternative solution $\mathbf{z}^{\prime}$ can lead to a non-identifiability result.
  • For case that there exists an weight $\lambda_{j,i}(\mathbf{u})+b$, we can always construct new latent variables $\mathbf{z^{\prime}}$ as: for all $k\neq i$, ${z^{\prime}}_{k}=z_{k}$, and $z^{\prime}_{i}=z_{i}-bz_{j}$. Given this, we can construct a matrix $\mathbf{M}$, so that $\mathbf{M}\mathbf{z^{\prime}}=\mathbf{z}$. It is clear the matrix $\mathbf{M}$ is invertible. In addition, all the weights of the matrix $\mathbf{M}$ are constant and thus do not depend on $\mathbf{u}$. Finally, we can construct a mapping from $\mathbf{z}^{\prime}$ to $\mathbf{x}$, e.g., $\mathbf{f}(\mathbf{M}\mathbf{z}^{\prime})$, which do not depend on $\mathbf{u}$, and can create the same data $\mathbf{x}$.
A.6 Implementation Details
A.6.1 Experiments for Figure 7
Data
For experimental results of Figure 7 the number of segments (e.g.,$\mathbf{u}$) $M$ is 10,20,30,40 and 50 respectively. For each segment, the number of latent causal variables is 2 and the sample size is 1000. We consider the following structural causal model
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\qquad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),$ (58)
${z_{1}}:=n_{1},\qquad\qquad\qquad\qquad\quad\ \ {z_{2}}:=\lambda_{1,2}(\mathbf{u}){z_{1}}+n_{2},$ (59)
where we sample the mean ${\beta}(\mathbf{u})_{i,1}$ and variance ${\beta}(\mathbf{u})_{i,2}$ from uniform distributions $[-2,2]$ and $[0.01,3]$, respectively. We sample the weights $\lambda_{1,2}(\mathbf{u})$ from uniform distributions $[0.1,2]$. We sample latent variable z according to Eqs. 58 and 59, and then mix them using a 2-layer multi-layer perceptron (MLP) to generate observed $\mathbf{x}$.
Network and Optimization
For all methods, we used a encoder, e.g. 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions, and decoder, e.g. 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions. We also use 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions. For optimization, we use Adam optimizer with learning rate $31r=1e-$ 3. For hyper-parameters, we set $\beta=4$ for the $\beta$-VAE. For CausalVAE, we set the hyper-parameters as suggested in the paper (Yang et al., 2021).
A.6.2 Experiments for Figures 6 and 8
Data
For experimental results of Figures 6 and 8, the number of segments is $30$ and sample size is 1000, while the number (e.g., dimension) of latent causal variables is 2,3,4,5 respectively. For 2-dimensional latent causal variables, the setting is the same as experiments for Figure 7 (a). For 3-dimensional case, we consider the following structural causal model:
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\quad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),\quad n_{3}:={\cal N}({\beta}(\mathbf{u})_{3,1},{\beta}(\mathbf{u})_{3,2}),$ (60)
${z_{1}}:=n_{1},\quad\qquad\qquad\qquad\quad\ \ {z_{2}}:=\lambda_{1,2}(\mathbf{u}){z_{1}}+n_{2},\qquad\quad\ {z_{3}}:=\lambda_{2,3}(\mathbf{u}){z_{2}}+n_{3}.$ (61)
For 4-dimensional case, we consider the following structural causal model:
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\qquad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),$ (62)
$n_{3}:={\cal N}({\beta}(\mathbf{u})_{3,1},{\beta}(\mathbf{u})_{3,2}),\qquad n_{4}:={\cal N}({\beta}(\mathbf{u})_{4,1},{\beta}(\mathbf{u})_{4,2}),$ (63)
${z_{1}}:=n_{1},\quad\qquad\qquad\qquad\quad\quad\ \ {z_{2}}:=\lambda_{1,2}(\mathbf{u}){z_{1}}+n_{2},$ (64)
${z_{3}}:=\lambda_{2,3}(\mathbf{u}){z_{2}}+n_{3},\qquad\ \qquad{z_{4}}:=\lambda_{2,4}(\mathbf{u}){z_{2}}+n_{4}.$ (65)
For 5-dimensional case, we consider the following structural causal model:
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\quad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),$ (66)
$n_{3}:={\cal N}({\beta}(\mathbf{u})_{3,1},{\beta}(\mathbf{u})_{3,2}),\quad n_{4}:={\cal N}({\beta}(\mathbf{u})_{4,1},{\beta}(\mathbf{u})_{4,2}),\quad n_{5}:={\cal N}({\beta}(\mathbf{u})_{5,1},{\beta}(\mathbf{u})_{5,2}),$ (67)
${z_{1}}:=n_{1},\quad\quad\qquad\qquad\quad\quad\ \ {z_{2}}:=\lambda_{1,2}(\mathbf{u}){z_{1}}+n_{2},$ (68)
${z_{3}}:=\lambda_{2,3}(\mathbf{u}){z_{2}}+n_{3},\qquad\ \quad{z_{4}}:=\lambda_{2,4}(\mathbf{u}){z_{2}}+n_{4},\qquad\quad\ {z_{5}}:=\lambda_{4,5}(\mathbf{u}){z_{4}}+n_{5}.$ (69)
Again, for all these Eqs 60-69, we sample the mean ${\beta}(\mathbf{u})_{i,1}$ and variance ${\beta}(\mathbf{u})_{i,2}$ from uniform distributions $[-2,2]$ and $[0.01,3]$, respectively. We sample the weights $\lambda_{i,j}(\mathbf{u})$ from uniform distributions $[0.1,2]$. We sample latent variable z according to these equations, and then mix them using a 2-layer multi-layer perceptron (MLP) to generate observed $\mathbf{x}$.
Network and Optimization
The network architecture, optimization and hyper-parameters are the same as used for experiments of Figure 7 (a), except for the 5-dimensional case where we use 40 hidden nodes for each linear layer.
A.6.3 Experiments for Table 1
For experimental results of Table 1, the number of segments is $30$ and the sample size is 1000, and the number (e.g., dimension) of latent causal variables is 3. For the case of the matching assumptions, we consider the structural causal model as mentioned in Eqs. (60)-(61). For the cases of uniform, Laplace and Gamma noise, we replace the original Gaussian noise by using uniform noise, Laplace and Gamma noise, respectively, while the remaining parts are the same as Eqs. (60)-(61). For the case of unchanged weights, we replace the weights in Eqs. (60)-(61) by unchanged weights across $\mathbf{u}$, which are sampled from an uniform distribution $[0.1,2]$. For the case of violated assumption (v), we replace the weights in Eqs. (60)-(61) by $\lambda_{i,j}(\mathbf{u})+b$, where $b$ is a constant from an uniform distribution $[0.1,2]$. The network architecture, optimization, and hyper-parameters are the same as used for experiments of Figure 7 (a).
A.7 Detailed Results on fMRI
Figure 21 shows the recovered latent six signals (Blue) and the true ones (Red) within one day by the proposed SuaVE.
Figure 21: Recovered latent six signals (Blue) and the true ones (Red) within one day by the proposed SuaVE.
The recovered latent fMRI variables within one day by iVAE are depicted by Figure 22.
Figure 22: Recovered latent six signals (Blue) and the true ones (Red) within one day by iVAE.
The recovered latent fMRI variables within one day by VAE are depicted by Figure 23.
Figure 23: Recovered latent six signals (Blue) and the true ones (Red) within one day by VAE.
The recovered latent fMRI variables within one day by $\beta$-VAE are depicted by Figure 24.
The recovered latent fMRI variables within one day by CausalVAE are depicted by Figure 25.
Figure 24: Recovered latent six signals (Blue) and the true ones (Red) within one day by $\beta$-VAE.
Figure 25: Recovered latent six signals (Blue) and the true ones (Red) within one day by CausalVAE.
Implementation Details: For fMRI data, again, we used the same network architecture for encoder (e.g. 3-layer fully connected network with 30 hidden nodes for each layer) and decoder (e.g. 3-layer fully connected network with 30 hidden nodes for each layer) parts in all these models. For prior model in the proposed SuaVE and iVAE, we use 3-layer fully connected network with 30 hidden nodes for each layer. We assign an 3-layer fully connected network with 30 nodes to generate the weights to model the relations among latent causal variables in the proposed SuaVE. For hyper-parameters, we set $\beta=4,25,50$ for $\beta$-VAE. For CausalVAE, we use the hyper-parameters setting as recommended by the authors (Yang et al., 2021), for generating observed $\mathbf{x}$, we use an invertible 2-layer multi-layer perceptron on the fMRI with different 5 random seeds.